### v1

# Energy Matching: Unifying Flow Matching and Energy-Based Models for Generative Modeling

Abstract
Generative models often map noise to data by matching flows or scores, but these approaches become cumbersome for incorporating partial observations or additional priors. Inspired by recent advances in Wasserstein gradient flows, we propose Energy Matching, a framework that unifies flow-based approaches with the flexibility of energy-based models (EBMs). Far from the data manifold, samples move along curl-free, optimal transport paths from noise to data. As they approach the data manifold, an entropic energy term guides the system into a Boltzmann equilibrium distribution, explicitly capturing the underlying likelihood structure of the data. We parameterize this dynamic with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems. Our method substantially outperforms existing EBMs on CIFAR-10 generation (FID $3.97$ compared to $8.61$), while retaining the simulation-free training of transport-based approaches away from the data manifold. Additionally, we exploit the flexibility of our method and introduce an interaction energy for diverse mode exploration. Our approach focuses on learning a static scalar potential energy—without time conditioning, auxiliary generators, or additional networks—marking a significant departure from recent EBM methods. We believe that this simplified framework significantly advances EBMs capabilities and paves the way for their broader adoption in generative modeling across diverse domains.
1 Introduction
[…]
Figure 1: Shown are trajectories (green lines) of samples traveling from a noise distribution (black dots; here, a Gaussian mixture model) to a data distribution (blue dots; here, two moons as in (Tong et al., 2023)) under four different methods: Action Matching (Neklyudov et al., 2023), Flow Matching (OT-CFM) (Tong et al., 2023), EBMs trained via contrastive divergence (Hinton, 2002), and our proposed Energy Matching. We highlight several individual trajectories in red to illustrate their distinct behaviors. Both Action Matching and Flow Matching learn time-dependent transports and are not trained for traversing the data manifold. Conversely, EBMs and Energy Matching are driven by time-independent fields that can be iterated indefinitely, allowing trajectories to navigate across modes. While samples from EBMs often require additional steps to equilibrate (see, e.g., the visible mode collapses that slow down sampling from the data manifold), Energy Matching directs samples toward the data distribution in “straight” paths, without hindering the exploration of the data manifold.
Many generative models learn a mapping from a simple, easy-to-sample distribution (e.g., a Gaussian) to the data distribution. They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023) — or via iterative noising-and-denoising schemes, as in diffusion models (Ho et al., 2020; Song et al., 2021). Besides being highly effective at sample generation, diffusion and flow-based models have also been used as priors to regularize ill-posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024). However, these models do not capture the unconditional data score explicitly, and instead model the score of smoothed manifolds at different noise levels. The measurement likelihood, on the other hand, is not tractable on these smoothed manifolds. As a result, existing approaches repeatedly shuttle between noised and data distributions, leading to crude approximations of complex, intractable terms Daras et al. (2024). For instance, DPS (Chung et al., 2023) approximates an intractable integral using a single sample. More recently, D-Flow (Ben-Hamu et al., 2024) optimizes the initial noise by differentiating through the simulated trajectory. To the best of our knowledge, these models lack a direct way to navigate the data manifold in search of the optimal solution without repeatedly transitioning between noised and data distributions.
EBMs (Hopfield, 1982; Hinton, 2002; LeCun et al., 2006) provide an alternative approach for approximating the data distribution by learning a scalar-valued function $E(x)$ that specifies an unnormalized density $p(x)\propto\exp\left(-E(x)\right)$. Rather than explicitly mapping noise samples onto the data manifold, EBMs assign low energies to regions of high data concentration and high energy elsewhere. This defines a Boltzmann distribution, from which one can sample by, for instance, Langevin sampling. In doing so, EBMs explicitly retain the likelihood information in $E(x)$. This likelihood information can then be leveraged in conditional generation (e.g., to solve inverse problems), possibly together with additional priors by simply adding their energy terms (Du and Mordatch, 2019). Moreover, direct examination of local curvature on the data manifold—allows the computation of local intrinsic dimension (LID) (an important proxy for data complexity)—whereas diffusion models can only approximate such curvature in the proximity of noise samples.
Despite the theoretical elegance of using a single, time-independent scalar energy, practical EBMs have historically suffered from poor generation quality, falling short of the performance of diffusion or flow matching models. Traditional methods (Song and Kingma, 2021) for training EBMs, such as contrastive divergence via Markov chain Monte Carlo (MCMC) or local score-based approaches (Song and Ermon, 2019), often fail to adequately explore the energy landscape in high-dimensional spaces, leading to instabilities and mode collapse. Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks (Guo et al., 2023; Zhang et al., 2024), thereby requiring significantly higher parameter counts and training complexity. Although multi-network approaches can improve sample quality, they inflate parameter counts by redundantly learning the same features across multiple networks, compromising the shared statistical strength of a unified feature set and the flexibility of time-independent models.
{mdframed}[hidealllines=true,backgroundcolor=blue!5]
Contributions.
In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure 1. When samples lie far from the data manifold, they are transported efficiently toward the data. Once near the data manifold, the flow transitions into Langevin steps governed by an internal energy component, enabling precise exploration of the Boltzmann-like density well around the data distribution. This straightforward approach produces a time-independent scalar energy field whose gradient both accelerates sampling and shapes the final density well—via a contrastive objective that directly learns the score at the data manifold—yet remains efficient and stable to train. Empirically, our method significantly outperforms existing EBMs on CIFAR-10 generation (FID $3.97$ vs. $8.61$) and compares favorably to flow-matching and diffusion models, without auxiliary generators or time-dependent EBM ensembles.
Our proposed process complements the advantages of flow matching with an explicit likelihood modeling, enabling traversal of the data manifold without repeatedly shuffling between noise and data distributions. This simplifies both inverse problem solving and controlled generations under a prior. In addition, to encourage diverse exploration of the data distribution, we showcase how repulsive interaction energies can be easily and effectively incorporated. Finally, we also showcase how analyzing the learned energy reveals insight on the LID of the data with fewer approximations than diffusion models.
2 Energy matching
In this section, we show how a scalar potential $V(x)$ can simultaneously provide an optimal-transport-like flow from noise to data while also yielding a Boltzmann distribution that explicitly captures the unnormalized likelihood of the data.
The Jordan–Kinderlehrer–Otto (JKO) scheme.
The starting point of our approach is the JKO scheme (Jordan et al., 1998), which is the basis of the success of numerous recent generative models (Xu et al., 2023; Terpin et al., 2024; Choi et al., 2024). The JKO scheme describes the discrete-time evolution of a probability distribution $\rho_{t}$ along energy-minimizing trajectories in the Wasserstein space,
$\rho_{t+\Delta t}=\arg\min_{\rho}\underbrace{\frac{W_{2}^{2}\left(\rho,\rho_{t}\right)}{2\Delta t}}_{\text{Transport Cost}}+\underbrace{\int V_{\theta}(x)\mathrm{d}\rho(x)}_{\text{Potential Energy}}+\underbrace{\varepsilon(t)\int\rho(x)\log\rho(x)\mathrm{d}x}_{\text{Internal Energy (-\text{Entropy)}}}.$ (1)
Here, $\theta$ denotes the learnable parameters of the scalar potential $V_{\theta}(x)$, and $\varepsilon(t)$ is a temperature-like parameter tuning the entropic term. The transport cost is given by the Wasserstein distance,
$W_{2}^{2}(\rho,\rho_{t})=\min_{\gamma\in\Gamma(\rho,\rho_{t})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\norm{x-x_{t}}^{2}\mathrm{d}\gamma(x,x_{t}),$ (2)
where $\Gamma(\rho,\rho_{t})$ is the set of couplings between $\rho$ and $\rho_{t}$, i.e., the set of probability distributions on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\rho$ and $\rho_{t}$. Here, $d$ is the dimensionality of the data. Henceforth, we call OT coupling any $\gamma_{t}$ that yields the minimum in (2). When $\gamma_{t}=(\mathrm{id},T)_{\#}\rho$, i.e., it is the pushforward of the map $x\mapsto(x,T(x))$ for some function $T$, we say that $T$ is an OT map from $\rho$ to $\rho_{t}$.
Differently from most literature, we consider $\varepsilon(t)$ to be dependent on time and study the behavior of Equation 1 as $t\to\infty$. To fix the ideas, consider, for instance, a linear scheduling:
$\varepsilon(t)=\begin{cases}0,&0\leq t<\operatorname{\tau^{\ast}},\\ \varepsilon_{\max}\frac{t-\operatorname{\tau^{\ast}}}{1-\operatorname{\tau^{\ast}}},&\operatorname{\tau^{\ast}}\leq t\leq 1,\\ \varepsilon_{\max},&t\geq 1.\end{cases}$ (3)
First-order optimality conditions.
We follow the approach in Terpin et al. (2024) and study (1) at each time $t$ via its first-order optimality conditions (Lanzetti et al., 2024, 2025):
$\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)+\varepsilon(t)\nabla_{x}\log\rho_{t}(x)=0,\qquad(x,y)\in\mathrm{supp}(\gamma_{t})$ (4)
where $\gamma_{t}$ is an OT plan between the distributions $\rho_{t}$ and $\rho_{t+\Delta t}$ and $\mathrm{supp}(\gamma_{t})$ is the support of $\gamma_{t}$. That is, this condition has to hold for all pairs of points in the support of $\rho_{t}$ and $\rho_{t+\Delta t}$ that are coupled by OT. Intuitively, analyzing (4) provides us with two key insights:
  1. 1.
    For times $t<\tau^{*}$, $\varepsilon(t)=0$ and (4) becomes
    $\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)=0\qquad(x,y)\in\mathrm{supp}(\gamma_{t}).$ (5)
    That is, the system is in an OT, flow-like, regime.
  2. 2.
    Near the data manifold, which we aim at modeling with the equilibrium distribution $\rho_{\text{eq}}$ of (1), $\rho_{t}\approx\rho_{\text{eq}}$ and, thus, for $t\gg 1$, $x\approx y$ for all $(x,y)\in\mathrm{supp}(\gamma_{t})$. Then, we can simplify (4) as
    $\varepsilon_{\max}\nabla_{x}\log\rho_{\mathrm{eq}}(x)=\nabla_{x}V_{\theta}(x)\quad\Longrightarrow\quad\rho_{\mathrm{eq}}(x)\propto\exp\left(-\tfrac{V_{\theta}(x)}{\varepsilon_{\max}}\right).$
    Thus, the equilibrium distribution is described by an EBM, $\exp\left(-E(x)\right)$, with $E(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$.
Our approach in a nutshell.
Combining the two insights above, we propose a generative framework that combines OT and EBMs to learn a time-independent scalar potential $V_{\theta}(x)$ whose Boltzmann distribution,
$\rho_{\mathrm{eq}}(x)\propto\exp\left(-\frac{V_{\theta}(x)}{\varepsilon_{\max}}\right),$ (6)
matches $\rho_{\mathrm{data}}$. To transport samples efficiently from noise $\rho_{0}$ to $\rho_{\mathrm{eq}}\approx\rho_{\mathrm{data}}$, we use two regimes:
  • Away from the data manifold: $\varepsilon\approx 0$. The flow is deterministic and OT-like, allowing rapid movement across large distances in sample space.
  • Near the data manifold: $\varepsilon\approx\varepsilon_{\max}$. Samples diffuse into a stable Boltzmann distribution, properly covering all data modes.
By combining the long-range transport capability of flows with the local density modeling flexibility of EBMs, we achieve tractable sampling and explicit likelihood of samples $V_{\theta}(x)$; see Figure 1.
2.1 Training objectives
In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase. Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach equilibrium.
2.1.1 Flow-like objective $\operatorname{\mathcal{L}_{\mathrm{OT}}}$
We begin by constructing a global velocity field $\nabla_{x}V_{\theta}(x)$ that carries noise samples $\{x_{0}\}$ to data samples $\{x_{\mathrm{data}}\}$ with minimal detours. For this, we consider geodesics in the Wasserstein space (Ambrosio et al., 2008). Practically, we compute the OT coupling $\gamma^{*}$ between two uniform empirical probability distributions, one supported on a mini-batch $\{x_{\text{data}}^{(j)}\}_{j=1}^{B}$ of the data, and one supported on a set of noise samples with the same cardinality $B$. These samples are drawn from an easy-to-sample distribution; in our case, a Gaussian. Since the probability distributions are uniform and empirical with the same number of samples, a transport map $T$ is guaranteed to exist (Ambrosio et al., 2008).
Remark 2.1 (OT solver).
Depending on the approach used to compute the OT coupling, one may find an OT map or not. Similarly, if the set of noise samples has cardinality smaller or larger than $B$, the OT coupling will not be expressed by a map. In this case, one can adapt our algorithm by defining a threshold $\pi_{\mathrm{th}}$ and consider all the pairs
$\left(x_{0}^{(j)},x_{\mathrm{data}}^{(i)}\right)\quad\text{with}\quad\gamma^{\ast}(x_{0}^{(j)},x_{\mathrm{data}}^{(i)})>\pi_{\mathrm{th}}.$
In our experiments, we did not notice benefits in a sample size other than $B$, in line with the observations in (Tong et al., 2023), and we used the POT solver (Flamary et al., 2021).
Then, for each data point $x_{\mathrm{data}}^{(i)}$ we define the interpolation
$x_{t}^{(i)}=(1-t)T(x_{\mathrm{data}}^{(i)})+tx_{\mathrm{data}}^{(i)},$
which is a point along the geodesic (Ambrosio et al., 2008). The velocity of each $x_{t}^{(i)}$ is $x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $-\nabla_{x}V_{\theta}(x_{t}^{(i)})\approx x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$. For this, we define the loss:
$\operatorname{\mathcal{L}_{\mathrm{OT}}}=\mathbb{E}^{\begin{subarray}{c}x_{\mathrm{data}}\in\mathcal{D}\\ t\sim U(0,1)\end{subarray}}\left[\norm{\nabla_{x}V_{\theta}(x_{t}^{(i)})+x_{\mathrm{data}}^{(i)}-T({x^{(i)}_{\mathrm{data}}})}^{2}\right].$
This objective can be interpreted as a flow-matching formulation under the assumption that the velocity field is both time-independent and given by the gradient of a scalar potential, thereby imposing a curl-free condition. This aligns naturally with OT, which also yields a velocity field of zero curl—any rotational component would add unnecessary distance to the flow and thus inflate the transport cost without benefit. Our experimental evidence adds to the recent study by (Sun et al., 2025), in which the authors observed that time-independent velocity fields can, under certain conditions, outperform time-dependent noise-conditioned fields in sample generation.
Algorithm 1 Phase 1 (warm-up).
1:Initialize model parameters $\theta$
2:for iteration $0,1,\ldots$ do
3:    Sample $\{x_{\mathrm{data}}^{(j)}\}_{j=1}^{B}\sim\mathcal{D}$triangleright Mini-batch of data samples
4:    Sample $\{x_{0}^{(i)}\}_{i=1}^{B}\sim\mathcal{N}(0,I)$triangleright Random Gaussian samples
5:    $T\leftarrow\mathrm{OTsolver}(\{x_{0}^{(i)}\},\{x_{\mathrm{data}}^{(j)}\})$triangleright Compute OT map
6:    Sample $t\sim U(0,1)$
7:    $\operatorname{\mathcal{L}_{\mathrm{OT}}}(\theta)\leftarrow\sum_{i}\|\nabla_{x}V_{\theta}(x_{t}^{(i)})+x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})\|^{2}$triangleright Loss function
8:    $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\text{OT}}(\theta)$triangleright Gradient update with learning rate $\alpha$
9:end for
10:return $\theta$$\triangleright$ Trained $\theta$
2.1.2 Contrastive objective $\operatorname{\mathcal{L}_{\mathrm{CD}}}$
Near the data manifold, $V_{\theta}(x)$ is refined so that $\rho_{\mathrm{eq}}(x)\propto\exp\left(-V_{\theta}(x)/\varepsilon_{\max}\right)$ matches the data distribution. We adopt the contrastive divergence loss described in EBMs (Hinton, 2002),
$\operatorname{\mathcal{L}_{\mathrm{CD}}}=\mathbb{E}_{x\sim p_{\text{data}}}\left[\frac{V_{\theta}(x)}{\varepsilon_{\max}}\right]-\mathbb{E}_{\tilde{x}\sim\mathrm{sg}(p_{\mathrm{eq}})}\left[\frac{V_{\theta}(\tilde{x})}{\varepsilon_{\max}}\right],$
where $\tilde{x}$ are “negative” samples of the equilibrium distribution induced by $V_{\theta}$. We approximate these samples using an MCMC Langevin chain (Welling and Teh, 2011). We split the initialization for negative samples: half begin at real data, and half begin at the noise distribution. This way, $V_{\theta}(x)$ forms well-defined basins around high-density regions while also shaping regions away from the manifold, correcting the generation. The $\mathrm{sg}(\cdot)$ indicates a stop-gradient operator, which ensures gradients do not back-propagate through the sampling procedure.
2.1.3 Dual objective
To balance the flow-like loss $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ with the contrastive divergence loss $\operatorname{\mathcal{L}_{\mathrm{CD}}}$, we adopt the linear temperature schedule (3). We define a sampling time $\operatorname{\tau_{s}}$. In fact, although convergence to the equilibrium distribution is guaranteed only as $\operatorname{\tau_{s}}\to\infty$, we empirically observe that sample quality (measured with Fréchet inception distance (FID)) plateaus by $\operatorname{\tau_{s}}=3.0$ on CIFAR-10; see Section A.2. The time discretization determines the temperature samples $\varepsilon^{(m)}\vcentcolon=\varepsilon(m\Delta t)$. We define a dataset-specific hyperparameter $\lambda_{\mathrm{CD}}$ to scale $\operatorname{\mathcal{L}_{\mathrm{CD}}}$. The resulting algorithm is described in detail in Algorithm 1 and Algorithm 2. Since Algorithm 2 benefits from high-quality negatives, we begin with Algorithm 1 (and, thus, with $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ only) to ensure sufficient mixing of noise-initialized negatives. Section A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases. Hyper-parameters for each dataset are defined in Appendix C.
Algorithm 2 Phase 2 (main training).
1:$\theta\leftarrow\theta_{\text{pretrained}}$triangleright Initialize from pretrained model (Algorithm 1)
2:for iteration $0,1,\ldots$ do
3:    Sample mini-batch $\{x_{\mathrm{data}}^{(j)}\}_{j=1}^{B}\sim\mathcal{D}$triangleright Mini-batch of data samples
4:    $\operatorname{\mathcal{L}_{\mathrm{OT}}}\leftarrow$ Use lines 4–7 from Algorithm 1
5:    Initialize $\{x_{\mathrm{neg}}^{(0)}\}$ from noise and/or data $\triangleright$ Negative samples
6:    for $m=0,1,\ldots,M_{\text{Langevin}}$ do
7:         Sample $\eta\sim\mathcal{N}(0,I)$
8:         $x_{\mathrm{neg}}^{(m+1)}\leftarrow x_{\mathrm{neg}}^{(m)}-\Delta t\nabla_{x}V_{\mathrm{sg}(\theta)}(x_{\mathrm{neg}}^{(m)})+\sqrt{2\Delta t\varepsilon^{(m)}}\eta$triangleright Langevin dynamics given $\varepsilon(t)$
9:         if $x_{\mathrm{neg}}^{(0)}$ initialized from data then
10:             Use $\varepsilon_{\max}$ instead of $\varepsilon^{(m)}$triangleright Use maximum temperature if initialized at the data manifold
11:         end if
12:    end for
13:    $\mathcal{L}_{\mathrm{neg}}\leftarrow\mathbb{E}[V_{\theta}(x_{\mathrm{neg}}^{(M)})]$
14:    $\mathcal{L}_{\mathrm{pos}}\leftarrow\mathbb{E}_{\mathrm{data}}[V_{\theta}(x_{\mathrm{data}})]$$\triangleright$ Compute expected potential energy on the data
15:    $\operatorname{\mathcal{L}_{\mathrm{CD}}}\leftarrow\mathcal{L}_{\mathrm{pos}}-\mathcal{L}_{\mathrm{neg}}$
16:    $\mathcal{L}(\theta)\leftarrow\operatorname{\mathcal{L}_{\mathrm{OT}}}+\lambda_{\mathrm{CD}}\operatorname{\mathcal{L}_{\mathrm{CD}}}$
17:    Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta)$triangleright Gradient descent step
18:end for
19:return $\theta$triangleright Trained $\theta$
3 Applications
In this section, we demonstrate the effectiveness and versatility of our proposed Energy Matching approach across three applications: (i) unconditional generation (ii) inverse problems, and (iii) LID estimation. The model architecture and all the training details are reported in Appendix C.
3.1 Unconditional generation
We compare four classes of generative models: (1) Diffusion models, which deliver state-of-the-art quality but typically require many sampling steps; (2) Flow-based methods, which learn OT paths for more efficient sampling with fewer steps; (3) EBMs, which directly model the log-density as a scalar field, offering flexibility for inverse problems and constraints but sometimes at the expense of sample quality; and (4) Ensembles (Diffusion + one or many EBMs), which combine diffusion’s robust sampling with elements of EBM flexibility but can become large and complex to train. Our approach, Energy Matching, offers a simple (a single time-independent scalar field) yet powerful EBM-based framework. We evaluate on CIFAR-10, reporting FID in Table 1. Our method outperforms state-of-the-art EBMs reducing FID by $50\%$.
Table 1: FID score comparison for unconditional CIFAR-10 generation (lower is better). Unless otherwise specified, we use results for solvers that most closely match our setup (300 fixed-step Euler–Heun (Butcher, 2016)). indicates reproduced methods, while all other entries reflect the best reported results.
Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories
Method FID $\downarrow$ Method FID $\downarrow$
Ensembles: Diffusion + (one or many) EBMs Diffusion Models
Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.54
EGC (Guo et al., 2023) 5.36 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45
Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68
Energy-based Models Flow-based Models
ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07
CLEL-large (38M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35
Energy Matching (50M params, Ours) 3.97 OT-CFM (37M params) (Tong et al., 2023) 4.11
3.2 Inverse problems
In many practical applications, we are interested in recovering some data $x$ from noisy measurements $y$ generated by an operator $A$, $y=A(x)+w$, where $w$ is Gaussian noise. In this setting, the posterior distribution of $x$ given $y$ is
$p(x|y)\propto\underbrace{\exp\left(-\|y-A(x)\|^{2}\right)}_{\displaystyle\propto p(y|x)}\underbrace{\exp\left(-E_{\theta}(x)\right)}_{\displaystyle\propto p(x)},$ (7)
where $E_{\theta}(x)$ is an energy function which one can learn from the data, an EBM. Because we want to sample $x$ given a measurement $y$, this reconstruction task is often referred to as an inverse problem. By taking the negative logarithm of (7), the maximum a posteriori estimate of $x$ given $y$ can be obtained by minimizing
$U_{\theta}(x;y)=\frac{1}{\zeta^{2}}\|y-A(x)\|^{2}+E_{\theta}(x).$ (8)
Here, $\|y-Ax\|^{2}$ encodes the measurement fidelity with $\zeta$ controlling the balance between this fidelity term and the prior. We obtain the prior term $E_{\theta}(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$ by training $V_{\theta}(x)$ via Energy Matching. Samples from this posterior can be drawn by starting from a random sample $x^{(0)}\sim\mathcal{N}(0,I)$ and following a Langevin update.
Suppose now we want to recover two images from a masked image, while additionally encouraging diverse reconstructions. EBMs readily allow us to do so by introducing an interaction energy
$W(x_{1},x_{2})=\frac{\norm{B(x_{1}-x_{2})}^{2}}{\sigma^{2}},$
where $B$ is a matrix with ones in the region of interest and zeros elsewhere, thus focusing the diversity within that particular fragment of the image, and $\sigma$ is a hyper-parameter controlling the strength of the interaction. Specifically, we define
$p(x_{1},x_{2}|y)\propto p(x_{1}|y)\cdot p(x_{2}|y)\cdot\exp\left(W(x_{1},x_{2})\right),$
giving high probability to pairs $(x_{1},x_{2})$ that lie far apart in the specified region $B$. This encourages exploring the edges of the posterior rather than just its modes. With suitable $W$, samples shift towards rare events without needing many draws. To illustrate the interaction term’s advantages for diverse reconstruction, we apply our method to a CelebA (Liu et al., 2015) $64\times 64$ inpainting task. From a partially observed (masked) face, we aim to reconstruct two distinct high-fidelity completions:
[…]
On the left is the masked face. On the right are the reconstructions: the top pair without the interaction term and the bottom pair with it. The interaction term applies in the solid red square (where $B$ has ones, zeros elsewhere), and the measurement matrix $A$ appears in the dotted blue square (zeros inside, ones outside). The interaction yields a wider range of completions yet preserves fidelity. We detail the inverse problems generation algorithm in Algorithm 3.
Algorithm 3 Sampling for inverse problems (with optional interaction energy)
1:for $m=1$ to $M$ do
2:    Initialize $x_{m}^{(0)}\sim\mathcal{N}(0,I)$triangleright Start each chain from Gaussian prior
3:end for
4:$N\leftarrow\lfloor\operatorname{\tau_{s}}/\Delta t\rfloor$triangleright Number of Langevin steps for sampling time $\operatorname{\tau_{s}}$
5:for $n=0$ to $N-1$ do
6:    for $m=1$ to $M$ do $\triangleright$ Data fidelity + prior + interaction energy
7:         $U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}\leftarrow\frac{\varepsilon^{(n)}}{\zeta^{2}}\bigl{\lVert}y-A\bigl{(}x_{m}^{(n)}\bigr{)}\bigr{\rVert}^{2}+V_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\frac{\varepsilon^{(n)}}{\sigma^{2}}\sum_{k\neq m}W\bigl{(}x_{m}^{(n)},x_{k}^{(n)}\bigr{)}$
8:         $\eta\sim\mathcal{N}(0,I)$triangleright Gaussian noise for Langevin step
9:         $x_{m}^{(n+1)}\leftarrow x_{m}^{(n)}-\Delta t\,\nabla_{x}U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\sqrt{2\varepsilon^{(n)}\Delta t}\;\eta$triangleright Langevin step
10:    end for
11:end for
12:return $\{x_{m}^{(N)}\}_{m=1}^{M}$triangleright Final set of samples
3.3 Local intrinsic dimension estimation
Real-world datasets, despite displaying a high number of variables, they can often be represented by a lower-dimensional manifold—a concept referred to as the manifold hypothesis (Fefferman et al., 2016). The dimension of this manifold is called the intrinsic dimension. Estimating the LID (Vapnik, 1995) at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities. We defer the precise definition and additional details to Appendix B.
Diffusion-based approaches.
Recent work leverages pre-trained diffusion models to estimate the LID (Kamkari et al., 2024; Stanczuk et al., 2024) by examining the learned score function. However, because these models do not learn the score at $t=0$ (the data manifold), their estimates become unreliable there. As a result, current methods rely on approximations, for instance, by evaluating the score at small but nonzero times $t=t_{0}$, where the computations remain sufficiently reliable.
Hessian-based LID Estimation.
Unlike diffusion models, EBMs explicitly parametrize the relative data likelihood. This explicit parametrization enables efficient analysis of the curvature of the underlying data manifold – in this example, estimating the LID. To this end, we compute the Hessian matrix $\nabla_{x}^{2}V(x_{\mathrm{data}})$ at a given data point and perform its spectral decomposition. We define near-zero eigenvalues as those whose absolute values lie within a small threshold $\tau$ (in our experiments, we set $\tau=10$ for MNIST(Deng, 2012) and $\tau=2$ for CIFAR-10). The count of near-zero eigenvalues reflects the number of flat directions and thus reveals the local dimension. As shown in Table 2, the LID estimates we obtain exhibit consistently stronger correlations with PNG compression size [^1] PNG is a lossless compression scheme specialized for images and can be used when no LID ground truth is available Kamkari et al. (2024). (evaluated on 4096 images) using Spearman’s correlation. Figure 2 offers qualitative illustrations. Our EBM-based approach outperforms diffusion-based methods by relying on fewer approximations, since the computation is performed exactly on the data manifold rather than in its proximity.
Spearman’s correlation $\uparrow$ MNIST CIFAR-10
ESS (Johnsson et al., 2014) 0.444 0.326
FLIPD (Kamkari et al., 2024) 0.837 0.819
NB (Stanczuk et al., 2024) 0.864 0.894
Energy Matching (Ours) 0.895 0.903
Table 2: We summarize the Spearman’s correlation coefficients of LID estimates with PNG compression rate. Benchmarks results reported in (Kamkari et al., 2024).
Figure 2: Qualitative results for LID estimation using the Hessian spectrum of $V_{\theta}(x)$. Left: Spectrum for a low-LID image. Right: Spectrum for a high-LID image. The eigenvalues quantify curvature along principal directions (eigenvectors). A degenerate spectrum (many near-zero eigenvalues) indicates locally "flat" regions, revealing the dimension of the local data manifold (LID). Intuitively, higher image complexity often corresponds to a higher LID.
4 Conclusion and limitations
Contributions.
We introduced a generative framework, Energy Matching, that reconciles the advantages of EBMs and OT flow matching models for simulation-free likelihood estimation and efficient high-fidelity generation. Specifically, it:
  • Learns a static scalar potential energy whose gradient drives rapid high-fidelity sampling–surpassing state-of-the-art energy based models–while also forming a Boltzmann-like density well suitable for controlled generation. All without auxiliary generators.
  • Offers efficient sampling from target data distributions on par with the state-of-the-art, while learning the score at the data manifold with minimal trainable parameters overhead.
  • Offers a simulation-free, principled likelihood estimation framework for solving inverse problems—where additional priors can be easily introduced—and enables the estimation of a data point’s LID with fewer approximations than score-based methods.
Limitations.
First, our method requires the extra gradient computation with respect to input, which can increase GPU memory usage (e.g. 20-40%), in particular during training. Second, for very high-resolution datasets, computing the full Hessian spectrum ($O(N^{3})$) may be impractical; and partial spectrum methods like random projections or iterative solvers may be used instead.
Outlook.
In contrast to the widespread belief, we demonstrated that static curl-free methods for generative flows work and offer an exciting direction for future research. Our Energy Matching approach has the potential to yield novel insights into conditional generation and inverse problems for cancer research Weidner et al. (2024); Balcerak et al. (2024), molecules and proteins Wu et al. (2022); Bilodeau et al. (2022), and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.
Appendix A Additional details on Energy Matching
In this section, we provide additional studies and visualizations on our method.
A.1 Energy landscape during training
In Figure 3, we visualize how the potential $V_{\theta}(x)$ transitions from a flow-like regime, where the OT loss enforces nearly zero curvature away from the data manifold (a), to an EBM-like regime, where the curvature around the new data geometry (here, two moons) is adaptively increased to approximate $\log p_{\text{data}}(x)$ (b). This two-stage design yields a well-shaped landscape that is both efficient to sample (thanks to a mostly flat potential between clusters) and accurate for density estimation near the data modes. For comparison, (c) shows an EBM trained solely with contrastive divergence, exhibiting sharper but less globally consistent basins.
a) Phase 1 (warm-up):
$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
b) Phase 2 (main training):
$\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}}$
c) Reference EBM:
$\operatorname{\mathcal{L}_{\mathrm{CD}}}$
Figure 3: Visualization of the energy $V_{\theta}(x)$ landscapes driving the samples from eight Gaussians to two moons. See Figure 1 for the 2D perspective. (a) The OT flow loss enforces zero curvature in $V_{\theta}(x)$ along the trajectories to the target. (b) Around the 2 Moons, the curvature of $V_{\theta}(x)$ is adjusted to approximate $\log p_{\text{moons}}(x)\propto V_{\theta}(x)$ while remaining close to the pretrained landscape elsewhere. Combining these objectives yields a potential energy landscape that is both efficient for sampling and representative of the underlying target data distribution. (c) An EBM is shown for comparison, trained using contrastive divergence loss. Visible mode collapse that slows down the equilibration. Less regular landscape away from the data as it needs many simulations to explore it.
A.2 Ablation on the sampling time
Here, we provide ablation studies on CIFAR-10 unconditional generation. Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$. This procedure produces a stable Boltzmann distribution from which one can sample. In the case of $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$. The reason for this is that once the samples move close to the data manifold there is no Boltzmann-like potential well to keep them from drifting away. We report results for different $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation 3).
11112222333300222244446666888810101010Sampling time $\operatorname{\tau_{s}}$ vs sampling time $\operatorname{\tau_{s}}$$\operatorname{\tau^{\ast}}=0.9,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\tau^{\ast}}=1.0,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
Figure 4: CIFAR-10 unconditional generation FID vs. sampling time $\operatorname{\tau_{s}}$ when sampling from models trained under different scenarios: pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ and combined $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, with temperature regime switching parameter $\operatorname{\tau^{\ast}}\in\{0.9,1.0\}$. Lower FID indicates better generative quality. All results for Euler-Heun with $\Delta t=0.01$.
Appendix B Details on LID estimation
Definition.
To start, we need to introduce the concept of local mass, defined as
[…]
where $p(x)$ is the local density and $B(x_{\text{data}},r)$ is a ball of radius $r$ around $x_{\text{data}}$, i.e. $B(x_{\text{data}},r)=\{x\in\mathbb{R}^{d}:\|x-x_{\mathrm{data}}\|\leq r\}$. The LID is then given by:
$\mathrm{LID}(x_{\mathrm{data}})=d-\lim_{r\to 0}\frac{\log\bigl(M(r)\bigr{missing})}{\log(r)}.$
Intuitively, $M(r)$ measures how much probability mass is concentrated in a ball of radius $r$ around $x_{\mathrm{data}}$. As we shrink this ball, the growth rate of $M(r)$ in terms of $r$ reveals the local dimensional structure of the data.
Assumptions.
In the context of contrastive divergence, we assume that data points $x_{\mathrm{data}}$ lie in well-like regions [Hyvärinen, 2006], i.e.:
[…]
Conceptually, $V(x)$ can be thought of as an energy function; points where $\nabla V(x_{\mathrm{data}})=0$ are near local minima of this energy, and the Hessian $\nabla^{2}V(x_{\mathrm{data}})$ provides information about local curvature (see Figure 2 for a qualitative illustration).
Energy-based density.
We define an energy-based density
$p(x)\;\propto\;\exp\!\Bigl{(}-\tfrac{V(x)}{\varepsilon}\Bigr{)},$
where $\varepsilon$ is a temperature parameter. Near a data point $x_{\mathrm{data}}$ satisfying $\nabla_{x}V(x_{\mathrm{data}})=0$, we can approximate $V(x)$ by its second-order Taylor expansion:
$V(x)\approx V(x_{\mathrm{data}})+\frac{1}{2}(x-x_{\mathrm{data}})^{\top}\nabla_{x}^{2}V(x_{\mathrm{data}})(x-x_{\mathrm{data}}).$
Consequently, in view of the assumptions above,
$p(x)\propto\exp\left(-\frac{1}{2\varepsilon}(x-x_{\mathrm{data}})^{\top}\nabla_{x}^{2}V(x_{\mathrm{data}})(x-x_{\mathrm{data}})\right).$
Local mass derivation and the rank of the energy Hessian.
Substituting the local quadratic form of $p(x)$ near $x_{\mathrm{data}}$ into the definition of the local mass $M(r)$, we obtain:
[…]
For small $r$, the dominant contribution depends on the rank of the Hessian $\nabla_{x}^{2}V(x_{\mathrm{data}})$. Let $k=\mathrm{rank}(\nabla_{x}^{2}V(x_{\mathrm{data}}))$. Then, as $r\to 0$, one can show that $M(r)=Cr^{k}$, where $C$ does not depend on $r$. We take the logarithm on both sides and divide by $\log(r)$ to get
$\frac{\log(M(r))}{\log(r)}=\frac{\log(C)+k\log(r)}{\log(r)}=k+\frac{\log(C)}{\log(r)},$
and the second term vanishes as $r\to 0$. Hence,
$\mathrm{LID}(x_{\mathrm{data}})=d-k.$
Practical estimation.
In practice, the LID at a data point $x_{\mathrm{data}}$ can be estimated through the following procedure:
  1. 1.
    Train $V(x)$ with Energy Matching.
  2. 2.
    Compute the Hessian $H=\nabla_{x}^{2}V(x_{\mathrm{data}})$.
  3. 3.
    Perform an eigenvalue decomposition on $H$.
Then the estimated local data-manifold dimension corresponds to the number of directions with negligible curvature (smaller magnitude than some $\tau$).
Appendix C Training Details
Below, we detail the training configurations for CIFAR-10, CelebA, and MNIST. The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd.
CIFAR-10.
The architecture is shown in Figure 5. We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively static) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output. Hyperparameters are: $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 200k iterations using Algorithm 1 and then 25k more with Algorithm 2 on 4xA100. The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=2\times 10^{-4}.$
CelebA.
We scale the CIFAR-10 model by 2$\sim$ 2$\times$ 2 ×; see Figure 6. We keep $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm 1 then 25k with Algorithm 2 on 4xA100. The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=2\times 10^{-5}$.
MNIST.
We train on a single RTX6000 directly with Algorithm 2 for 300k iterations using $\lambda_{\mathrm{CD}}=0.25$, $\Delta t=0.05$, and $M_{\text{Langevin}}=50$. The architecture consists of a two-stage CNN followed by a two-layer MLP, comprising a total of 2M parameters.
Input(3$\times$32$\times$32)UNet(37M params)Transformer(PatchEmbed+ViT, 12M params)ScalarPotential
Hyperparameter Value
Image size 3$\times$32$\times$32
Base channels (UNet) 128
ResBlocks 2
channel_mult [1, 2, 2, 2]
Attention resolution 16
Attention heads (UNet) 4
Head channels (UNet) 64
Dropout 0.1
Transformer (ViT) Head
Patch size 4
Embedding dim 384
Transformer layers 8
Transformer heads 4
Output scale 1000.0
Figure 5: Diagram of our UNet+Transformer EBM for CIFAR-10. A UNet (37M params) processes a 3$\times$32$\times$32 image; its output is fed into a Transformer head (PatchEmbed + 8-layer ViT, 12M params) that produces a scalar potential. Here we employ the identical UNet architecture as in [Tong et al., 2023].
Input(3$\times$64$\times$64)UNet(83M params)Transformer(PatchEmbed+ViT, 25M params)ScalarPotential
Hyperparameter Value
Image size 3$\times$64$\times$64
Base channels (UNet) 128
ResBlocks 2
channel_mult [1, 2, 3, 4]
Attention resolution 16
Attention heads (UNet) 4
Head channels (UNet) 64
Dropout 0.1
Transformer (ViT) Head
Patch size 4
Embedding dim 512
Transformer layers 8
Transformer heads 8
Output scale 1000.0
Figure 6: Diagram of our UNet+Transformer EBM for CelebA. A UNet (83M params) processes a 3$\times$64$\times$64 image; its output is fed into a Transformer head (PatchEmbed+8-layer ViT, 25M params) that produces a scalar potential.

### v3

@@
 Abstract
-Generative models often map noise to data by matching flows or scores, but these approaches become cumbersome for incorporating partial observations or additional priors.
-Inspired by recent advances in Wasserstein gradient flows, we propose Energy Matching, a framework that unifies flow-based approaches with the flexibility of energy-based models (EBMs).
+The most widely used generative models map noise and data distributions by matching flows or scores.
+However, they struggle to incorporate partial observations and additional priors—something energy-based models (EBMs) handle elegantly by simply adding corresponding scalar energy terms.
+We address this issue by proposing Energy Matching, a framework that endows flow-based approaches with the flexibility of EBMs.
 Far from the data manifold, samples move along curl-free, optimal transport paths from noise to data.
@@
 We parameterize this dynamic with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems.
-Our method substantially outperforms existing EBMs on CIFAR-10 generation (FID $3.97$ compared to $8.61$), while retaining the simulation-free training of transport-based approaches away from the data manifold.
-Additionally, we exploit the flexibility of our method and introduce an interaction energy for diverse mode exploration.
-Our approach focuses on learning a static scalar potential energy—without time conditioning, auxiliary generators, or additional networks—marking a significant departure from recent EBM methods.
-We believe that this simplified framework significantly advances EBMs capabilities and paves the way for their broader adoption in generative modeling across diverse domains.
+Our method substantially outperforms existing EBMs on CIFAR-10 and ImageNet generation in terms of fidelity, while retaining simulation-free training of transport-based approaches away from the data manifold.
+Furthermore, we leverage the method’s flexibility to introduce an interaction energy that supports diverse mode exploration, which we demonstrate in a controlled protein-generation setting.
+Our approach focuses on learning a scalar potential energy—without time-conditioning, auxiliary generators, or additional networks—which marks a significant departure from recent EBM methods.
+We believe that this simplified framework significantly advances EBMs capabilities and paves the way for their wider adoption in generative modeling across diverse domains.
 1 Introduction
 […]
-Figure 1: Shown are trajectories (green lines) of samples traveling from a noise distribution (black dots; here, a Gaussian mixture model) to a data distribution (blue dots; here, two moons as in (Tong et al., 2023)) under four different methods: Action Matching (Neklyudov et al., 2023), Flow Matching (OT-CFM) (Tong et al., 2023), EBMs trained via contrastive divergence (Hinton, 2002), and our proposed Energy Matching.
+Figure 1: Trajectories (green lines) of samples traveling from a noise distribution (black dots; here, a Gaussian mixture model) to a data distribution (blue dots; here, two moons as in (Tong et al., 2023)) under four different methods: Action Matching (Neklyudov et al., 2023), Flow Matching (OT-CFM) (Tong et al., 2023), EBMs trained via contrastive divergence (Hinton, 2002), and our proposed Energy Matching.
 We highlight several individual trajectories in red to illustrate their distinct behaviors.
@@
 While samples from EBMs often require additional steps to equilibrate (see, e.g., the visible mode collapses that slow down sampling from the data manifold), Energy Matching directs samples toward the data distribution in “straight” paths, without hindering the exploration of the data manifold.
-Many generative models learn a mapping from a simple, easy-to-sample distribution (e.g., a Gaussian) to the data distribution.
-They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023) — or via iterative noising-and-denoising schemes, as in diffusion models (Ho et al., 2020; Song et al., 2021).
-Besides being highly effective at sample generation, diffusion and flow-based models have also been used as priors to regularize ill-posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024).
-However, these models do not capture the unconditional data score explicitly, and instead model the score of smoothed manifolds at different noise levels.
-The measurement likelihood, on the other hand, is not tractable on these smoothed manifolds.
+Many generative models learn to map from a simple, easy-to-sample distribution, such as a Gaussian, to a desired data distribution.
+They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023) - or through iterative noising and denoising schemes, such as in diffusion models (Ho et al., 2020; Song et al., 2021).
+In addition to being highly effective in sample generation, diffusion- and flow-based models have also been used as priors to regularize poorly posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024).
+However, these models do not explicitly capture the unconditional data score and instead model the score of smoothed manifolds at different noise levels.
+The measurement likelihood, on the other hand, is not tractable on these noised manifolds.
 As a result, existing approaches repeatedly shuttle between noised and data distributions, leading to crude approximations of complex, intractable terms Daras et al.
 (2024).
-For instance, DPS (Chung et al., 2023) approximates an intractable integral using a single sample.
-More recently, D-Flow (Ben-Hamu et al., 2024) optimizes the initial noise by differentiating through the simulated trajectory.
+For example, DPS (Chung et al., 2023) approximates an intractable integral using a single sample.
+More recently, D-Flow (Ben-Hamu et al., 2024) optimizes initial noise by differentiating through the simulated trajectory.
 To the best of our knowledge, these models lack a direct way to navigate the data manifold in search of the optimal solution without repeatedly transitioning between noised and data distributions.
@@
 Rather than explicitly mapping noise samples onto the data manifold, EBMs assign low energies to regions of high data concentration and high energy elsewhere.
-This defines a Boltzmann distribution, from which one can sample by, for instance, Langevin sampling.
+This defines a Boltzmann distribution from which one can sample, for example, via Langevin sampling.
 In doing so, EBMs explicitly retain the likelihood information in $E(x)$.
-This likelihood information can then be leveraged in conditional generation (e.g., to solve inverse problems), possibly together with additional priors by simply adding their energy terms (Du and Mordatch, 2019).
+This likelihood information can then be used in conditional generation (e.g., to solve inverse problems), possibly together with additional priors simply by adding their energy terms (Du and Mordatch, 2019).
 Moreover, direct examination of local curvature on the data manifold—allows the computation of local intrinsic dimension (LID) (an important proxy for data complexity)—whereas diffusion models can only approximate such curvature in the proximity of noise samples.
@@
 Traditional methods (Song and Kingma, 2021) for training EBMs, such as contrastive divergence via Markov chain Monte Carlo (MCMC) or local score-based approaches (Song and Ermon, 2019), often fail to adequately explore the energy landscape in high-dimensional spaces, leading to instabilities and mode collapse.
-Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks (Guo et al., 2023; Zhang et al., 2024), thereby requiring significantly higher parameter counts and training complexity.
-Although multi-network approaches can improve sample quality, they inflate parameter counts by redundantly learning the same features across multiple networks, compromising the shared statistical strength of a unified feature set and the flexibility of time-independent models.
+Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks trained in cooperation (Guo et al., 2023; Zhang et al., 2024; Yoon et al., 2024), thereby requiring significantly higher parameter counts and training complexity.
 {mdframed}[hidealllines=true,backgroundcolor=blue!5]
@@
 In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure 1.
-When samples lie far from the data manifold, they are transported efficiently toward the data.
+When samples lie far from the data manifold, they are efficiently transported toward the data.
 Once near the data manifold, the flow transitions into Langevin steps governed by an internal energy component, enabling precise exploration of the Boltzmann-like density well around the data distribution.
 This straightforward approach produces a time-independent scalar energy field whose gradient both accelerates sampling and shapes the final density well—via a contrastive objective that directly learns the score at the data manifold—yet remains efficient and stable to train.
-Empirically, our method significantly outperforms existing EBMs on CIFAR-10 generation (FID $3.97$ vs.
-$8.61$) and compares favorably to flow-matching and diffusion models, without auxiliary generators or time-dependent EBM ensembles.
+Empirically, our method significantly outperforms existing EBMs on both CIFAR-10 and ImageNet generation in terms of fidelity, and compares favorably to flow-matching and diffusion models—without auxiliary generators or time-dependent EBM ensembles.
 Our proposed process complements the advantages of flow matching with an explicit likelihood modeling, enabling traversal of the data manifold without repeatedly shuffling between noise and data distributions.
 This simplifies both inverse problem solving and controlled generations under a prior.
-In addition, to encourage diverse exploration of the data distribution, we showcase how repulsive interaction energies can be easily and effectively incorporated.
+In addition, to encourage diverse exploration of the data distribution, we showcase how repulsive interaction energies can be easily and effectively incorporated, with an application to conditional protein generation.
 Finally, we also showcase how analyzing the learned energy reveals insight on the LID of the data with fewer approximations than diffusion models.
+[^1] The code will be publicly released upon acceptance.
 2 Energy matching
@@
 2.1 Training objectives
-In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase.
-Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach equilibrium.
+In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$, see Section 2.1.1), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase.
+Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach equilibrium (see Section 2.1.2).
 2.1.1 Flow-like objective $\operatorname{\mathcal{L}_{\mathrm{OT}}}$
@@
 Similarly, if the set of noise samples has cardinality smaller or larger than $B$, the OT coupling will not be expressed by a map.
-In this case, one can adapt our algorithm by defining a threshold $\pi_{\mathrm{th}}$ and consider all the pairs
-$\left(x_{0}^{(j)},x_{\mathrm{data}}^{(i)}\right)\quad\text{with}\quad\gamma^{\ast}(x_{0}^{(j)},x_{\mathrm{data}}^{(i)})>\pi_{\mathrm{th}}.$
-In our experiments, we did not notice benefits in a sample size other than $B$, in line with the observations in (Tong et al., 2023), and we used the POT solver (Flamary et al., 2021).
-Then, for each data point $x_{\mathrm{data}}^{(i)}$ we define the interpolation
-$x_{t}^{(i)}=(1-t)T(x_{\mathrm{data}}^{(i)})+tx_{\mathrm{data}}^{(i)},$
-which is a point along the geodesic (Ambrosio et al., 2008).
+In this case, one can adapt our algorithm by defining a threshold $\pi_{\mathrm{th}}$ and consider all the pairs $\left(x_{0}^{(j)},x_{\mathrm{data}}^{(i)}\right)$ with $\gamma^{\ast}(x_{0}^{(j)},x_{\mathrm{data}}^{(i)})>\pi_{\mathrm{th}}$.
+In our experiments, we used the POT solver (Flamary et al., 2021) and we did not notice benefits in a sample size other than $B$, in line with the observations in (Tong et al., 2023).
+Then, for each data point $x_{\mathrm{data}}^{(i)}$ we define the interpolation $x_{t}^{(i)}=(1-t)T(x_{\mathrm{data}}^{(i)})+tx_{\mathrm{data}}^{(i)}$, which is a point along the geodesic (Ambrosio et al., 2008).
 The velocity of each $x_{t}^{(i)}$ is $x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $-\nabla_{x}V_{\theta}(x_{t}^{(i)})\approx x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$.
@@
 Section A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases.
-Hyper-parameters for each dataset are defined in Appendix C.
+Hyper-parameters for each dataset are defined in Appendix D.
 Algorithm 2 Phase 2 (main training).
-1:$\theta\leftarrow\theta_{\text{pretrained}}$triangleright Initialize from pretrained model (Algorithm 1)
+1:$\theta\leftarrow\theta_{\text{pretrained}}$triangleright Initialize from the warmed-up model (Algorithm 1)
 2:for iteration $0,1,\ldots$ do
@@
 In this section, we demonstrate the effectiveness and versatility of our proposed Energy Matching approach across three applications: (i) unconditional generation (ii) inverse problems, and (iii) LID estimation.
-The model architecture and all the training details are reported in Appendix C.
+The model architecture and all the training details are reported in Appendix D.
 3.1 Unconditional generation
-We compare four classes of generative models: (1) Diffusion models, which deliver state-of-the-art quality but typically require many sampling steps; (2) Flow-based methods, which learn OT paths for more efficient sampling with fewer steps; (3) EBMs, which directly model the log-density as a scalar field, offering flexibility for inverse problems and constraints but sometimes at the expense of sample quality; and (4) Ensembles (Diffusion + one or many EBMs), which combine diffusion’s robust sampling with elements of EBM flexibility but can become large and complex to train.
+We compare four classes of generative models: (1) Diffusion models, which deliver state-of-the-art quality but typically require many sampling steps; (2) Flow-based methods, which learn OT paths for more efficient sampling with fewer steps; (3) EBMs, which directly model the log-density as a scalar field, offering flexibility for inverse problems and constraints but sometimes at the expense of sample quality; and (4) Ensembles (Diffusion with one or many EBMs), which combine diffusion’s robust sampling with elements of EBM flexibility but can become large and complex to train.
 Our approach, Energy Matching, offers a simple (a single time-independent scalar field) yet powerful EBM-based framework.
-We evaluate on CIFAR-10, reporting FID in Table 1.
-Our method outperforms state-of-the-art EBMs reducing FID by $50\%$.
-Table 1: FID score comparison for unconditional CIFAR-10 generation (lower is better).
-Unless otherwise specified, we use results for solvers that most closely match our setup (300 fixed-step Euler–Heun (Butcher, 2016)).
-indicates reproduced methods, while all other entries reflect the best reported results.
+We evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009) and ImageNet32x32 (Deng et al., 2009; Chrabaszcz et al., 2017) datasets, reporting FID scores in Table 1 and Table 2.
+Our method surpasses state-of-the-art EBMs, achieving a $50\%$ reduction in FID.
+[…]
+.
+Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories Ensembles: Diffusion + (one or many) EBMs Diffusion Models Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.54 EGC (Guo et al., 2023) 5.36 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 3.45 Cooperative DRL (40M params) (Zhu et al., 2024) 4.31 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45 Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68 Energy-based Models Flow-based Models ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07 CLEL-large (32M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35 Energy Matching (50M params, Ours) 3.97 OT-CFM (37M params) (Tong et al., 2023) 4.11
+[…]
 Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories
-Method FID $\downarrow$ Method FID $\downarrow$
 Ensembles: Diffusion + (one or many) EBMs Diffusion Models
-Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.54
-EGC (Guo et al., 2023) 5.36 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45
-Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68
+Cooperative DRL (40M params) (Zhu et al., 2024) 9.35 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 8.42
 Energy-based Models Flow-based Models
-ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07
-CLEL-large (38M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35
-Energy Matching (50M params, Ours) 3.97 OT-CFM (37M params) (Tong et al., 2023) 4.11
+ImprovedCD (Du et al., 2021) 32.48 Flow-matching (Lipman et al., 2023) (196M params) 5.02
+CLEL-base (Lee et al., 2023) (7M params) 22.16
+CLEL-large (Lee et al., 2023) (32M params) 15.47
+Energy Matching (50M params, Ours) 7.67
 3.2 Inverse problems
-In many practical applications, we are interested in recovering some data $x$ from noisy measurements $y$ generated by an operator $A$, $y=A(x)+w$, where $w$ is Gaussian noise.
+In many practical applications, we are interested in recovering some data $x$ from noisy measurements $y$ generated by an operator $A$, $y=A(x)+w$, where $w\sim\mathcal{N}\!\bigl{(}0,\sqrt{2}\,\zeta I\bigr{)}$.
 In this setting, the posterior distribution of $x$ given $y$ is
-$p(x|y)\propto\underbrace{\exp\left(-\|y-A(x)\|^{2}\right)}_{\displaystyle\propto p(y|x)}\underbrace{\exp\left(-E_{\theta}(x)\right)}_{\displaystyle\propto p(x)},$ (7)
+$p(x|y)\propto\underbrace{\exp\left(-\frac{1}{\zeta^{2}}\|y-A(x)\|^{2}\right)}_{\displaystyle\propto p(y|x)}\underbrace{\exp\left(-E_{\theta}(x)\right)}_{\displaystyle\propto p(x)},$ (7)
 where $E_{\theta}(x)$ is an energy function which one can learn from the data, an EBM.
 Because we want to sample $x$ given a measurement $y$, this reconstruction task is often referred to as an inverse problem.
-By taking the negative logarithm of (7), the maximum a posteriori estimate of $x$ given $y$ can be obtained by minimizing
-$U_{\theta}(x;y)=\frac{1}{\zeta^{2}}\|y-A(x)\|^{2}+E_{\theta}(x).$ (8)
 Here, $\|y-Ax\|^{2}$ encodes the measurement fidelity with $\zeta$ controlling the balance between this fidelity term and the prior.
@@
 Samples from this posterior can be drawn by starting from a random sample $x^{(0)}\sim\mathcal{N}(0,I)$ and following a Langevin update.
-Suppose now we want to recover two images from a masked image, while additionally encouraging diverse reconstructions. EBMs readily allow us to do so by introducing an interaction energy
-$W(x_{1},x_{2})=\frac{\norm{B(x_{1}-x_{2})}^{2}}{\sigma^{2}},$
-where $B$ is a matrix with ones in the region of interest and zeros elsewhere, thus focusing the diversity within that particular fragment of the image, and $\sigma$ is a hyper-parameter controlling the strength of the interaction.
-Specifically, we define
-$p(x_{1},x_{2}|y)\propto p(x_{1}|y)\cdot p(x_{2}|y)\cdot\exp\left(W(x_{1},x_{2})\right),$
-giving high probability to pairs $(x_{1},x_{2})$ that lie far apart in the specified region $B$.
-This encourages exploring the edges of the posterior rather than just its modes.
-With suitable $W$, samples shift towards rare events without needing many draws.
+We detail the algorithm for generating solutions to inverse problems in Algorithm 3 of Appendix B.
+We demonstrate our model’s capabilities qualitatively through a controlled inpainting task and quantitatively via a protein inverse design benchmark.
+Controlled inpainting.
+Suppose we want to recover two images from a masked image while encouraging diverse reconstructions.
+EBMs allow this by introducing an additional interaction energy, $W(x_{1},x_{2})=-\frac{\|B(x_{1}-x_{2})\|^{2}}{\sigma^{2}}$, where $B$ has ones in the region of interest (focusing diversity there) and zeros elsewhere, and $\sigma$ is a hyper-parameter controlling the interaction’s strength.
+Specifically, we define $p(x_{1},x_{2}\mid y)\propto p(x_{1}\mid y)\,p(x_{2}\mid y)\,\exp\bigl(-W(x_{1},x_{2})\bigr{missing})$, which gives high probability to pairs $(x_{1},x_{2})$ that lie far apart in the specified region $B$.
+This encourages exploring the edges of the posterior rather than just its modes, and with suitable $W$, samples shift toward rare events without needing many draws.
 To illustrate the interaction term’s advantages for diverse reconstruction, we apply our method to a CelebA (Liu et al., 2015) $64\times 64$ inpainting task.
-From a partially observed (masked) face, we aim to reconstruct two distinct high-fidelity completions:
+As shown in Figure 2, we start from a partially observed (masked) face and aim to reconstruct two distinct high-fidelity completions.
 […]
+Figure 2: Controlled inpainting for diverse reconstructions.
 On the left is the masked face.
-On the right are the reconstructions: the top pair without the interaction term and the bottom pair with it.
-The interaction term applies in the solid red square (where $B$ has ones, zeros elsewhere), and the measurement matrix $A$ appears in the dotted blue square (zeros inside, ones outside).
-The interaction yields a wider range of completions yet preserves fidelity.
-We detail the inverse problems generation algorithm in Algorithm 3.
-Algorithm 3 Sampling for inverse problems (with optional interaction energy)
-1:for $m=1$ to $M$ do
-2:    Initialize $x_{m}^{(0)}\sim\mathcal{N}(0,I)$triangleright Start each chain from Gaussian prior
-3:end for
-4:$N\leftarrow\lfloor\operatorname{\tau_{s}}/\Delta t\rfloor$triangleright Number of Langevin steps for sampling time $\operatorname{\tau_{s}}$
-5:for $n=0$ to $N-1$ do
-6:    for $m=1$ to $M$ do $\triangleright$ Data fidelity + prior + interaction energy
-7:         $U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}\leftarrow\frac{\varepsilon^{(n)}}{\zeta^{2}}\bigl{\lVert}y-A\bigl{(}x_{m}^{(n)}\bigr{)}\bigr{\rVert}^{2}+V_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\frac{\varepsilon^{(n)}}{\sigma^{2}}\sum_{k\neq m}W\bigl{(}x_{m}^{(n)},x_{k}^{(n)}\bigr{)}$
-8:         $\eta\sim\mathcal{N}(0,I)$triangleright Gaussian noise for Langevin step
-9:         $x_{m}^{(n+1)}\leftarrow x_{m}^{(n)}-\Delta t\,\nabla_{x}U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\sqrt{2\varepsilon^{(n)}\Delta t}\;\eta$triangleright Langevin step
-10:    end for
-11:end for
-12:return $\{x_{m}^{(N)}\}_{m=1}^{M}$triangleright Final set of samples
+On the right are two reconstructions: the top pair without the interaction term and the bottom pair with it.
+The interaction term applies in the solid red square (where $B$ has ones), and the measurement matrix $A$ is the dotted blue square (zeros inside, ones outside).
+By encouraging $x_{1}$ and $x_{2}$ to differ in the target region, the interaction yields a wider range of completions while preserving fidelity.
+Protein inverse design.
+In Figure 3, we demonstrate our method’s performance on the inverse design problem of generating Adeno-Associated Virus (AAV) capsid protein segments (Bryant et al., 2021).
+Given a desired functional property (fitness)—here defined as the predicted viral packaging efficiency normalized between 0 and 1—the goal is to design novel protein sequences satisfying this target condition.
+Beyond achieving high fitness, practical inverse design requires generating diverse candidate sequences to ensure robustness in the downstream experimental validation (Jain et al., 2022).
+We evaluate on two benchmark splits (medium and hard), which correspond to subsets of the original AAV dataset differing in baseline fitness distributions and required mutational distance from known high-performing variants (Kirjner et al., 2024).
+Leveraging the latent-space representation of VLGPO (Bogensperger et al., 2025), we employ our Energy-Matching Langevin sampler with an inference-time tunable repulsion term, allowing explicit control over the diversity of the designed proteins.
+This enables a flexible trade-off between fitness and diversity, resulting in high fitness scores alongside substantially improved sequence diversity.
+See Section B.2 for experimental details and dataset descriptions.
+[…]
+Figure 3: Fitness–diversity trade-off for protein inverse design on the AAV Medium (left) and Hard (right) benchmarks.
+We compare our Energy Matching method (blue), with diversity explicitly controlled by a repulsion strength parameter ($\propto\frac{1}{\sigma^{2}}$), against leading flow-based (purple), score-based (orange), and other non-likelihood methods (black).
+Fitness measures how well generated sequences satisfy the target property (predicted viral packaging efficiency), while diversity quantifies the average Levenshtein distance between sequences in each generated batch.
 3.3 Local intrinsic dimension estimation
-Real-world datasets, despite displaying a high number of variables, they can often be represented by a lower-dimensional manifold—a concept referred to as the manifold hypothesis (Fefferman et al., 2016).
-The dimension of this manifold is called the intrinsic dimension.
-Estimating the LID (Vapnik, 1995) at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities.
-We defer the precise definition and additional details to Appendix B.
+Real-world datasets, despite displaying a high number of variables, can often be represented by lower-dimensional manifolds—a concept referred to as the manifold hypothesis (Fefferman et al., 2016).
+The dimension of such a manifold is called the intrinsic dimension.
+Estimating the LID at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities.
+We defer the precise definition to Appendix C.
 Diffusion-based approaches.
@@
 The count of near-zero eigenvalues reflects the number of flat directions and thus reveals the local dimension.
-As shown in Table 2, the LID estimates we obtain exhibit consistently stronger correlations with PNG compression size [^1] PNG is a lossless compression scheme specialized for images and can be used when no LID ground truth is available Kamkari et al.
+As shown in Table 3, the LID estimates we obtain exhibit stronger correlations with PNG compression size [^2] PNG is a lossless compression scheme specialized for images and can provide useful guidance when no LID ground truth is available Kamkari et al.
 (2024).
@@
 Figure 4 offers qualitative illustrations.
-Our EBM-based approach outperforms diffusion-based methods by relying on fewer approximations, since the computation is performed exactly on the data manifold rather than in its proximity.
+Our EBM-based approach compares favorably to diffusion-based methods, as it relies on fewer approximations by performing computations exactly on the data manifold rather than merely in its vicinity.
 Spearman’s correlation $\uparrow$ MNIST CIFAR-10
@@
 NB (Stanczuk et al., 2024) 0.864 0.894
-Energy Matching (Ours) 0.895 0.903
-Table 2: We summarize the Spearman’s correlation coefficients of LID estimates with PNG compression rate. Benchmarks results reported in (Kamkari et al., 2024).
+Energy Matching (Ours) 0.895 0.901
+Table 3: Spearman’s correlation coefficients of LID estimates with PNG compression rate. Benchmarks results reported in (Kamkari et al., 2024).
 Figure 4: Qualitative results for LID estimation using the Hessian spectrum of $V_{\theta}(x)$.
@@
 The eigenvalues quantify curvature along principal directions (eigenvectors).
-A degenerate spectrum (many near-zero eigenvalues) indicates locally "flat" regions, revealing the dimension of the local data manifold (LID).
+A degenerate spectrum (many near-zero eigenvalues, marked in red) indicates locally "flat" regions, revealing the LID.
 Intuitively, higher image complexity often corresponds to a higher LID.
@@
 Specifically, it:
-  • Learns a static scalar potential energy whose gradient drives rapid high-fidelity sampling–surpassing state-of-the-art energy based models–while also forming a Boltzmann-like density well suitable for controlled generation.
+  • Learns a time-independent scalar potential energy whose gradient drives rapid high-fidelity sampling–surpassing state-of-the-art energy-based models–while also forming a Boltzmann-like density well suitable for controlled generation.
 All without auxiliary generators.
-  • Offers efficient sampling from target data distributions on par with the state-of-the-art, while learning the score at the data manifold with minimal trainable parameters overhead.
+  • Offers efficient sampling from target data distributions on par with the state-of-the-art, while learning the score at the data manifold with manageable trainable parameters overhead.
   • Offers a simulation-free, principled likelihood estimation framework for solving inverse problems—where additional priors can be easily introduced—and enables the estimation of a data point’s LID with fewer approximations than score-based methods.
@@
 Outlook.
-In contrast to the widespread belief, we demonstrated that static curl-free methods for generative flows work and offer an exciting direction for future research.
-Our Energy Matching approach has the potential to yield novel insights into conditional generation and inverse problems for cancer research Weidner et al.
+In contrast to the widespread belief, we demonstrated that time-independent curl-free methods for generative flows work and offer an exciting direction for future research.
+Our Energy Matching approach has the potential to yield novel insights into controlled generation and inverse problems for cancer research Weidner et al.
 (2024); Balcerak et al.
@@
 (2022); Bilodeau et al.
-(2022), and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.
+(2022), computational fluid dynamics Gao et al.
+(2024); Shysheya et al.
+(2024); Molinaro et al.
+(2024) and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.
 Appendix A Additional details on Energy Matching
@@
 Here, we provide ablation studies on CIFAR-10 unconditional generation.
-Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$.
-This procedure produces a stable Boltzmann distribution from which one can sample.
-In the case of $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$.
-The reason for this is that once the samples move close to the data manifold there is no Boltzmann-like potential well to keep them from drifting away.
-We report results for different $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation 3).
+Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, producing a stable Boltzmann distribution from which one can sample.
+Figure 6 illustrates the FID as a function of sampling time $\operatorname{\tau_{s}}$ for models trained under these different regimes.
+In the case of pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$; this occurs because, once the samples move close to the data manifold, there is no Boltzmann-like potential well to keep them from drifting away.
+Because the fidelity slope near the data manifold is steep with respect to sampling time, methods lacking explicit time-conditioning can easily overshoot or undershoot, significantly impacting fidelity.
+This behavior might explain why some models degrade in performance when made time-independent, as recently reported by Sun et al.
+[2025].
+In Figure 6 we also report results for different values of the temperature-switching parameter $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation 3).
 11112222333300222244446666888810101010Sampling time $\operatorname{\tau_{s}}$ vs sampling time $\operatorname{\tau_{s}}$$\operatorname{\tau^{\ast}}=0.9,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\tau^{\ast}}=1.0,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
@@
 All results for Euler-Heun with $\Delta t=0.01$.
-Appendix B Details on LID estimation
+Appendix B Posterior sampling
+This section provides additional algorithmic details and experimental settings for inverse problem sampling tasks presented in the main text.
+B.1 Algorithm description
+The detailed sampling procedure for inverse problems, optionally incorporating an interaction energy term to explicitly control the diversity among generated samples, is provided in Algorithm 3.
+The Langevin dynamics alternates between evaluating a composite potential $U_{\theta}(x)$, composed of a data fidelity term, a learned energy-based prior $V_{\theta}$, and, optionally, an interaction energy $W(x,x^{\prime})$.
+This interaction term can, for instance, encourage diversity among parallel chains.
+Samples are initialized independently from the Gaussian prior and iteratively refined through gradient-based updates.
+Specific hyperparameters, including the discretization step size $\Delta t$, total sampling time $T$, and the interaction energy strength, are detailed in Appendix D.
+Algorithm 3 Sampling for inverse problems (with optional interaction energy)
+1:for $m=1$ to $M$ do
+2:    Initialize $x_{m}^{(0)}\sim\mathcal{N}(0,I)$triangleright Start each chain from Gaussian prior
+3:end for
+4:$N\leftarrow\lfloor\operatorname{\tau_{s}}/\Delta t\rfloor$triangleright Number of Langevin steps for sampling time $\operatorname{\tau_{s}}$
+5:for $n=0$ to $N-1$ do
+6:    for $m=1$ to $M$ do $\triangleright$ Data fidelity + prior + interaction energy
+7:         $U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}\leftarrow\frac{\varepsilon^{(n)}}{\zeta^{2}}\bigl{\lVert}y-A\bigl{(}x_{m}^{(n)}\bigr{)}\bigr{\rVert}^{2}+V_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}-\frac{\varepsilon^{(n)}}{\sigma^{2}}\sum_{k\neq m}W\bigl{(}x_{m}^{(n)},x_{k}^{(n)}\bigr{)}$
+8:         $\eta\sim\mathcal{N}(0,I)$triangleright Gaussian noise for Langevin step
+9:         $x_{m}^{(n+1)}\leftarrow x_{m}^{(n)}-\Delta t\,\nabla_{x}U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\sqrt{2\varepsilon^{(n)}\Delta t}\;\eta$triangleright Langevin step
+10:    end for
+11:end for
+12:return $\{x_{m}^{(N)}\}_{m=1}^{M}$triangleright Final set of samples
+B.2 AAV inverse design protein generation
+We consider protein fitness optimization for the adeno-associated virus (AAV) in the medium and hard data regime as proposed by [Kirjner et al., 2024].
+The setup for the latent encoding as well as the backbone architecture for the learned prior in the latent space follows [Bogensperger et al., 2025].
+Conditional sampling uses classifier guidance via learned predictor networks $g_{\phi}$ or $\tilde{g}_{\phi}$ to guide samples toward high-fitness regions.
+The CNN-based networks predicting fitness for AAV sequences are directly sourced from [Kirjner et al., 2024] and are trained only on the limited training data for the medium and hard setup, respectively.
+For training we follow Algorithm 1 and Algorithm 2.
+We then sample 128 sequences using Algorithm 3 (same batch size for all baselines), the hyperparameters are detailed in Appendix D.
+Generated sequences are evaluated for fitness using the learned oracle from Kirjner et al.
+[2024].
+Further, they are assessed for diversity within the generated sequences and novelty relative to the training set [Jain et al., 2022].
+As shown in Table 4, our method achieves state-of-the-art fitness while boosting diversity, and incorporating interaction energy in Algorithm 3 further increases diversity with only a minimal impact on fitness.
+Table 4: AAV optimization results.
+For VLGPO (flow-based) and Energy Matching, medium difficulty uses $g_{\phi}$, hard difficulty uses $\tilde{g}_{\phi}$.
+Metrics (Fitness↑, Diversity↑, Novelty↑).
+Reported uncertainty of Fitness is expressed as standard deviation.
+AAV medium AAV hard
+Method Fitness↑ Diversity↑ Novelty↑ Fitness↑ Diversity↑ Novelty↑
+Learning Unnormalized Data Likelihood
+Energy-based Models
+Energy Matching (Ours) 0.59 (0.0) 5.86 5.0 0.61 (0.0) 4.77 6.7
+Energy Matching (+repulsion) (Ours) 0.58 (0.0) 6.22 5.0 0.60 (0.0) 5.22 6.6
+Learning Transport/Score Along Noised Trajectories
+Flow-based Models
+VLGPO [Bogensperger et al., 2025] 0.58 (0.0) 5.58 5.0 0.61 (0.0) 4.29 6.2
+Diffusion Models
+gg-dWJS [Ikram et al., 2024] 0.48 (0.0) 9.48 4.2 0.33 (0.0) 14.3 5.3
+Other Methods
+LatProtRL [Lee et al., 2024] 0.57 (0.0) 3.00 5.0 0.57 (0.0) 3.00 5.0
+GGS [Kirjner et al., 2024] 0.51 (0.0) 4.00 5.4 0.60 (0.0) 4.50 7.0
+AdaLead [Sinai et al., 2020] 0.46 (0.0) 8.50 2.8 0.40 (0.0) 8.53 3.4
+CbAS [Brookes et al., 2019] 0.43 (0.0) 12.70 7.2 0.36 (0.0) 14.4 8.6
+GWG [Grathwohl et al., 2021] 0.43 (0.1) 6.60 7.7 0.33 (0.0) 12.0 12.2
+GFN-AL [Jain et al., 2022] 0.20 (0.1) 9.60 19.4 0.10 (0.1) 11.6 19.6
+Appendix C Details on LID estimation
 Definition.
@@
 Then the estimated local data-manifold dimension corresponds to the number of directions with negligible curvature (smaller magnitude than some $\tau$).
-Appendix C Training Details
-Below, we detail the training configurations for CIFAR-10, CelebA, and MNIST. The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd.
+Appendix D Training details
+Below, we detail the training configurations for CIFAR-10, ImageNet 32x32, CelebA, MNIST, and AAV.
+The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd [Ansel et al., 2024].
+We optimize all models with the Adam optimizer Kingma and Ba [2014], and for CIFAR-10, ImageNet 32x32, and CelebA we maintain an exponential moving average (EMA) of the weights.
+While we specifically adopt (i) a one-sided trimmed mean of negative sample energies and (ii) clamping of the contrastive loss for stability, any commonly used EBM technique (e.g., persistent contrastive divergence [Tieleman, 2008], replay buffers, multi-scale negative sampling) could be readily employed.
+In our approach, we introduce two hyperparameters, $\alpha$ and $\beta$, to control these stabilizing techniques:
+[…]
+$\beta$ $=\text{clamp threshold for $\mathcal{L}_{\mathrm{CD}}$ (i.e., we clamp $\mathcal{L}_{\mathrm{CD}}$ to be $\geq-\beta$).}$.
 CIFAR-10.
 The architecture is shown in Figure 7.
-We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively static) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output.
+We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively time-independent) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output.
 Hyperparameters are: $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$.
-We train for 200k iterations using Algorithm 1 and then 25k more with Algorithm 2 on 4xA100.
-The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=2\times 10^{-4}.$
+We train for 200k iterations using Algorithm 1 with EMA $0.9999$ and then 10k more with Algorithm 2 and EMA $0.999$ on 4xA100.
+The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=8\times 10^{-3},\alpha=0.1,\beta=-0.05$.
+ImageNet 32x32.
+The architecture is shown in Figure 7 (same as for CIFAR-10).
+Hyperparameters are: $\tau_{s}=1.5$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$.
+We train for 600k iterations using Algorithm 1 with EMA $0.9999$ and then 7k more with Algorithm 2 and EMA $0.999$ on 7xA100.
+The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=2\times 10^{-4}.$ $\alpha=0.2,\beta=-0.01$.
 CelebA.
 We scale the CIFAR-10 model by 2$\sim$ 2$\times$ 2 ×; see Figure 8.
-We keep $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm 1 then 25k with Algorithm 2 on 4xA100.
-The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=2\times 10^{-5}$.
+We keep $\tau_{s}=2.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm 1 with EMA $0.9999$ then 7k with Algorithm 2 and EMA $0.999$ on 4xA100.
+The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
 MNIST.
-We train on a single RTX6000 directly with Algorithm 2 for 300k iterations using $\lambda_{\mathrm{CD}}=0.25$, $\Delta t=0.05$, and $M_{\text{Langevin}}=50$.
+We train on a single RTX6000 directly with Algorithm 2 for 300k iterations using $\Delta t=0.05$, $M_{\text{Langevin}}=50$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=0.25$.
 The architecture consists of a two-stage CNN followed by a two-layer MLP, comprising a total of 2M parameters.
+AAV.
+We adopt the one-dimensional CNN architecture as used in [Bogensperger et al., 2025], summing the final-layer activations to obtain the potential.
+We train for 10k iterations using Algorithm 1 and for 1k iterations using Algorithm 2 on a single GeForce RTX 4090.
+The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
+For Algorithm 3 we use $\tau_{s}=1.7$ for AAV medium and $\tau_{s}=1.3$ for AAV hard, $\tau^{*}=0.9$, $\zeta=0.01$ for AAV medium and $\zeta=0.009$ for AAV hard and $\Delta t=0.01$.
+We set the target fitness to $y=1$ to aim for the maximum fitness in the generated sequences.
 Input(3$\times$32$\times$32)UNet(37M params)Transformer(PatchEmbed+ViT, 12M params)ScalarPotential
@@
 Output scale 1000.0
-Figure 5: Diagram of our UNet+Transformer EBM for CIFAR-10.
+Figure 7: Diagram of our UNet+Transformer EBM for CIFAR-10 and ImageNet 32x32.
 A UNet (37M params) processes a 3$\times$32$\times$32 image; its output is fed into a Transformer head (PatchEmbed + 8-layer ViT, 12M params) that produces a scalar potential.
-Here we employ the identical UNet architecture as in [Tong et al., 2023].
+Here we employ the identical UNet architecture as in [Tong et al., 2023], but with the time parameter fixed at $t=0$ to render the model time-independent.
 Input(3$\times$64$\times$64)UNet(83M params)Transformer(PatchEmbed+ViT, 25M params)ScalarPotential

### v5

@@
 Abstract
-The most widely used generative models map noise and data distributions by matching flows or scores.
-However, they struggle to incorporate partial observations and additional priors—something energy-based models (EBMs) handle elegantly by simply adding corresponding scalar energy terms.
-We address this issue by proposing Energy Matching, a framework that endows flow-based approaches with the flexibility of EBMs.
-Far from the data manifold, samples move along curl-free, optimal transport paths from noise to data.
+Current state-of-the-art generative models map noise to data distributions by matching flows or scores.
+A key limitation of these models is their inability to readily integrate available partial observations and additional priors.
+In contrast, energy-based models (EBMs) address this by incorporating corresponding scalar energy terms.
+Here, we propose Energy Matching, a framework that endows flow-based approaches with the flexibility of EBMs.
+Far from the data manifold, samples move from noise to data along irrotational, optimal transport paths.
 As they approach the data manifold, an entropic energy term guides the system into a Boltzmann equilibrium distribution, explicitly capturing the underlying likelihood structure of the data.
-We parameterize this dynamic with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems.
-Our method substantially outperforms existing EBMs on CIFAR-10 and ImageNet generation in terms of fidelity, while retaining simulation-free training of transport-based approaches away from the data manifold.
-Furthermore, we leverage the method’s flexibility to introduce an interaction energy that supports diverse mode exploration, which we demonstrate in a controlled protein-generation setting.
-Our approach focuses on learning a scalar potential energy—without time-conditioning, auxiliary generators, or additional networks—which marks a significant departure from recent EBM methods.
-We believe that this simplified framework significantly advances EBMs capabilities and paves the way for their wider adoption in generative modeling across diverse domains.
+We parameterize these dynamics with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems.
+The present method substantially outperforms existing EBMs on CIFAR-10 and ImageNet generation in terms of fidelity, while retaining simulation-free training of transport-based approaches away from the data manifold.
+Furthermore, we leverage the flexibility of the method to introduce an interaction energy that supports the exploration of diverse modes, which we demonstrate in a controlled protein generation setting.
+This approach learns a scalar potential energy, without time conditioning, auxiliary generators, or additional networks, marking a significant departure from recent EBM methods.
+We believe this simplified yet rigorous formulation significantly advances EBMs capabilities and paves the way for their wider adoption in generative modeling in diverse domains.
 1 Introduction
@@
 While samples from EBMs often require additional steps to equilibrate (see, e.g., the visible mode collapses that slow down sampling from the data manifold), Energy Matching directs samples toward the data distribution in “straight” paths, without hindering the exploration of the data manifold.
-Many generative models learn to map from a simple, easy-to-sample distribution, such as a Gaussian, to a desired data distribution.
-They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023) - or through iterative noising and denoising schemes, such as in diffusion models (Ho et al., 2020; Song et al., 2021).
+Generative models learn to map from a simple, easy-to-sample distribution, such as a Gaussian, to a desired data distribution.
+They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Albergo and Vanden-Eijnden, 2023)—or through iterative noising and denoising schemes, such as in diffusion models (Ho et al., 2020; Song et al., 2021).
 In addition to being highly effective in sample generation, diffusion- and flow-based models have also been used as priors to regularize poorly posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024).
@@
 Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks trained in cooperation (Guo et al., 2023; Zhang et al., 2024; Yoon et al., 2024), thereby requiring significantly higher parameter counts and training complexity.
-{mdframed}[hidealllines=true,backgroundcolor=blue!5]
 Contributions.
-In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure 1.
+In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure˜1.
 When samples lie far from the data manifold, they are efficiently transported toward the data.
@@
 Finally, we also showcase how analyzing the learned energy reveals insight on the LID of the data with fewer approximations than diffusion models.
-[^1] The code will be publicly released upon acceptance.
+[^1] Code repository: https://github.com/m1balcerak/EnergyMatching
 2 Energy matching
-In this section, we show how a scalar potential $V(x)$ can simultaneously provide an optimal-transport-like flow from noise to data while also yielding a Boltzmann distribution that explicitly captures the unnormalized likelihood of the data.
+In this section, we show how a scalar potential $V(x)$ can simultaneously provide an optimal-transport-like flow from noise to data while also yielding a Boltzmann distribution that explicitly captures the unnormalized log‑likelihood of the data.
 The Jordan–Kinderlehrer–Otto (JKO) scheme.
@@
 The transport cost is given by the Wasserstein distance,
-$W_{2}^{2}(\rho,\rho_{t})=\min_{\gamma\in\Gamma(\rho,\rho_{t})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\norm{x-x_{t}}^{2}\mathrm{d}\gamma(x,x_{t}),$ (2)
+$W_{2}^{2}(\rho,\rho_{t})=\min_{\gamma\in\Gamma(\rho,\rho_{t})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\norm{x-y}^{2}\mathrm{d}\gamma(x,y),$ (2)
 where $\Gamma(\rho,\rho_{t})$ is the set of couplings between $\rho$ and $\rho_{t}$, i.e., the set of probability distributions on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\rho$ and $\rho_{t}$.
@@
 When $\gamma_{t}=(\mathrm{id},T)_{\#}\rho$, i.e., it is the pushforward of the map $x\mapsto(x,T(x))$ for some function $T$, we say that $T$ is an OT map from $\rho$ to $\rho_{t}$.
-Differently from most literature, we consider $\varepsilon(t)$ to be dependent on time and study the behavior of Equation 1 as $t\to\infty$.
+Differently from most literature, we consider $\varepsilon(t)$ to be dependent on time and study the behavior of Equation˜1 as $t\to\infty$.
 To fix the ideas, consider, for instance, a linear scheduling:
-$\varepsilon(t)=\begin{cases}0,&0\leq t<\operatorname{\tau^{\ast}},\\ \varepsilon_{\max}\frac{t-\operatorname{\tau^{\ast}}}{1-\operatorname{\tau^{\ast}}},&\operatorname{\tau^{\ast}}\leq t\leq 1,\\ \varepsilon_{\max},&t\geq 1.\end{cases}$ (3)
+$\varepsilon(t)=\begin{cases}0,&0\leq t<\operatorname{\tau^{\ast}},\\ \varepsilon_{\max}\frac{t-\operatorname{\tau^{\ast}}}{1-\operatorname{\tau^{\ast}}},&\operatorname{\tau^{\ast}}\leq t<1,\\ \varepsilon_{\max},&t\geq 1.\end{cases}$ (3)
 First-order optimality conditions.
-We follow the approach in Terpin et al. (2024) and study (1) at each time $t$ via its first-order optimality conditions (Lanzetti et al., 2024, 2025):
-$\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)+\varepsilon(t)\nabla_{x}\log\rho_{t}(x)=0,\qquad(x,y)\in\mathrm{supp}(\gamma_{t})$ (4)
-where $\gamma_{t}$ is an OT plan between the distributions $\rho_{t}$ and $\rho_{t+\Delta t}$ and $\mathrm{supp}(\gamma_{t})$ is the support of $\gamma_{t}$.
-That is, this condition has to hold for all pairs of points in the support of $\rho_{t}$ and $\rho_{t+\Delta t}$ that are coupled by OT.
+Following Terpin et al.
+(2024), we analyze (1) at each time $t$ via its first-order optimality conditions (Lanzetti et al., 2024, 2025).
+These conditions characterize the properties of the desired solution and thus represent the optimization goal:
+$\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)+\varepsilon(t)\nabla_{x}\log\rho_{t+\Delta t}(x)=0,\qquad(x,y)\in\mathrm{supp}(\gamma_{t})$ (4)
+where $\gamma_{t}$ is an OT plan between the distributions $\rho_{t+\Delta t}$ and $\rho_{t}$ and $\mathrm{supp}(\gamma_{t})$ is the support of $\gamma_{t}$.
+That is, this condition has to hold for all pairs of points in the support of $\rho_{t+\Delta t}$ and $\rho_{t}$ that are coupled by OT.
 Intuitively, analyzing (4) provides us with two key insights:
   1. 1.
-    For times $t<\tau^{*}$, $\varepsilon(t)=0$ and (4) becomes
+    For times $t<\operatorname{\tau^{\ast}}$, $\varepsilon(t)=0$ and (4) becomes
     $\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)=0\qquad(x,y)\in\mathrm{supp}(\gamma_{t}).$ (5)
@@
   2. 2.
-    Near the data manifold, which we aim at modeling with the equilibrium distribution $\rho_{\text{eq}}$ of (1), $\rho_{t}\approx\rho_{\text{eq}}$ and, thus, for $t\gg 1$, $x\approx y$ for all $(x,y)\in\mathrm{supp}(\gamma_{t})$.
+    Near the data manifold, which we aim at modeling with the equilibrium distribution $\rho_{\text{eq}}$ of (1), $\rho_{t+\Delta t}\approx\rho_{\text{eq}}$ and, thus, for $t\gg 1$, $x\approx y$ for all $(x,y)\in\mathrm{supp}(\gamma_{t})$.
 Then, we can simplify (4) as
-    $\varepsilon_{\max}\nabla_{x}\log\rho_{\mathrm{eq}}(x)=\nabla_{x}V_{\theta}(x)\quad\Longrightarrow\quad\rho_{\mathrm{eq}}(x)\propto\exp\left(-\tfrac{V_{\theta}(x)}{\varepsilon_{\max}}\right).$
+    $\varepsilon_{\max}\nabla_{x}\log\rho_{\mathrm{eq}}(x)=-\nabla_{x}V_{\theta}(x)\quad\Longrightarrow\quad\rho_{\mathrm{eq}}(x)\propto\exp\left(-\tfrac{V_{\theta}(x)}{\varepsilon_{\max}}\right).$
     Thus, the equilibrium distribution is described by an EBM, $\exp\left(-E(x)\right)$, with $E(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$.
@@
   • Near the data manifold: $\varepsilon\approx\varepsilon_{\max}$. Samples diffuse into a stable Boltzmann distribution, properly covering all data modes.
-By combining the long-range transport capability of flows with the local density modeling flexibility of EBMs, we achieve tractable sampling and explicit likelihood of samples $V_{\theta}(x)$; see Figure 1.
+By combining the long-range transport capability of flows with the local density modeling flexibility of EBMs, we achieve tractable sampling and explicitly encode the unnormalized log-likelihood $V_{\theta}(x)/\varepsilon_{\max}$ of the underlying data distribution; see Figure˜1.
 2.1 Training objectives
-In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$, see Section 2.1.1), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase.
-Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach equilibrium (see Section 2.1.2).
+In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$, see Section˜2.1.1), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase.
+Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach the data manifold (i.e., the equilibrium distribution); see Section˜2.1.2.
 2.1.1 Flow-like objective $\operatorname{\mathcal{L}_{\mathrm{OT}}}$
@@
 For this, we consider geodesics in the Wasserstein space (Ambrosio et al., 2008).
-Practically, we compute the OT coupling $\gamma^{*}$ between two uniform empirical probability distributions, one supported on a mini-batch $\{x_{\text{data}}^{(j)}\}_{j=1}^{B}$ of the data, and one supported on a set of noise samples with the same cardinality $B$.
+Practically, we compute the OT coupling $\gamma^{*}$ between two uniform empirical probability distributions, one supported on a mini-batch of the data, and one supported on a set of noise samples with the same cardinality.
 These samples are drawn from an easy-to-sample distribution; in our case, a Gaussian.
@@
 Remark 2.1 (OT solver).
-Depending on the approach used to compute the OT coupling, one may find an OT map or not.
-Similarly, if the set of noise samples has cardinality smaller or larger than $B$, the OT coupling will not be expressed by a map.
-In this case, one can adapt our algorithm by defining a threshold $\pi_{\mathrm{th}}$ and consider all the pairs $\left(x_{0}^{(j)},x_{\mathrm{data}}^{(i)}\right)$ with $\gamma^{\ast}(x_{0}^{(j)},x_{\mathrm{data}}^{(i)})>\pi_{\mathrm{th}}$.
-In our experiments, we used the POT solver (Flamary et al., 2021) and we did not notice benefits in a sample size other than $B$, in line with the observations in (Tong et al., 2023).
-Then, for each data point $x_{\mathrm{data}}^{(i)}$ we define the interpolation $x_{t}^{(i)}=(1-t)T(x_{\mathrm{data}}^{(i)})+tx_{\mathrm{data}}^{(i)}$, which is a point along the geodesic (Ambrosio et al., 2008).
-The velocity of each $x_{t}^{(i)}$ is $x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $-\nabla_{x}V_{\theta}(x_{t}^{(i)})\approx x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$.
+Depending on the method used to compute the OT coupling, an explicit OT map may or may not be obtained.
+Similarly, if the number of noise samples differs from the mini-batch size $B$, the resulting OT coupling generally will not correspond to a map.
+In this case, one can adapt the algorithm by defining a threshold $\pi_{\mathrm{th}}$ and considering all pairs $(x_{\mathrm{data}},x_{0})$ for which the coupling value satisfies $\gamma^{\ast}(x_{\mathrm{data}},x_{0})>\pi_{\mathrm{th}}$.
+In our experiments, we used the POT solver (Flamary et al., 2021) and did not observe benefits from using a sample size different from $B$, consistent with previous approaches (Tong et al., 2023).
+Then, for each data point $x_{\mathrm{data}}$ we define the interpolation $x_{t}=(1-t)T(x_{\mathrm{data}})+tx_{\mathrm{data}}$, which is a point along the geodesic.
+The velocity of each $x_{t}$ is $x_{\mathrm{data}}-T(x_{\mathrm{data}})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $\nabla_{x}V_{\theta}(x_{t})\approx x_{\mathrm{data}}-T(x_{\mathrm{data}})$.
 For this, we define the loss:
-$\operatorname{\mathcal{L}_{\mathrm{OT}}}=\mathbb{E}^{\begin{subarray}{c}x_{\mathrm{data}}\in\mathcal{D}\\ t\sim U(0,1)\end{subarray}}\left[\norm{\nabla_{x}V_{\theta}(x_{t}^{(i)})+x_{\mathrm{data}}^{(i)}-T({x^{(i)}_{\mathrm{data}}})}^{2}\right].$
-This objective can be interpreted as a flow-matching formulation under the assumption that the velocity field is both time-independent and given by the gradient of a scalar potential, thereby imposing a curl-free condition.
-This aligns naturally with OT, which also yields a velocity field of zero curl—any rotational component would add unnecessary distance to the flow and thus inflate the transport cost without benefit.
+$\operatorname{\mathcal{L}_{\mathrm{OT}}}=\mathbb{E}{\begin{subarray}{c}x_{\mathrm{data}}\in\mathcal{D}\\ t\sim U(0,\operatorname{\tau^{\ast}})\end{subarray}}\left[\norm{\nabla_{x}V_{\theta}(x_{t})+x_{\mathrm{data}}-T({x_{\mathrm{data}}})}^{2}\right].$
+This objective can be interpreted as a flow-matching formulation under the assumption that the velocity field is both time-independent and given by the gradient of a scalar potential, thereby imposing an irrotational condition.
+This aligns naturally with OT, which also yields an irrotational velocity field—any rotational component would add unnecessary distance to the flow and thus inflate the transport cost without benefit.
 Our experimental evidence adds to the recent study by (Sun et al., 2025), in which the authors observed that time-independent velocity fields can, under certain conditions, outperform time-dependent noise-conditioned fields in sample generation.
@@
 1:Initialize model parameters $\theta$
-2:for iteration $0,1,\ldots$ do
-3:    Sample $\{x_{\mathrm{data}}^{(j)}\}_{j=1}^{B}\sim\mathcal{D}$triangleright Mini-batch of data samples
-4:    Sample $\{x_{0}^{(i)}\}_{i=1}^{B}\sim\mathcal{N}(0,I)$triangleright Random Gaussian samples
-5:    $T\leftarrow\mathrm{OTsolver}(\{x_{0}^{(i)}\},\{x_{\mathrm{data}}^{(j)}\})$triangleright Compute OT map
-6:    Sample $t\sim U(0,1)$
-7:    $\operatorname{\mathcal{L}_{\mathrm{OT}}}(\theta)\leftarrow\sum_{i}\|\nabla_{x}V_{\theta}(x_{t}^{(i)})+x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})\|^{2}$triangleright Loss function
-8:    $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\text{OT}}(\theta)$triangleright Gradient update with learning rate $\alpha$
-9:end for
-10:return $\theta$$\triangleright$ Trained $\theta$
+2:for iteration $n=0,1,\ldots$ do
+3:  Sample mini-batch $\{x_{\mathrm{data},b}\}_{b=1}^{B}\sim\mathcal{D}$ $\triangleright$ Data samples
+4:  Sample mini-batch $\{x_{0,b}\}_{b=1}^{B}\sim\mathcal{N}(0,I)$ $\triangleright$ Random Gaussian samples
+5:  $T\leftarrow\mathrm{OTsolver}(\{x_{\mathrm{data},b}\},\{x_{0,b}\})$ $\triangleright$ Compute OT map
+6:  Sample $\{t_{b}\}_{b=1}^{B}\sim U(0,\operatorname{\tau^{\ast}})$ $\triangleright$ Typically $\operatorname{\tau^{\ast}}=1$ for the warm-up
+7:  Set interpolations $x_{t_{b}}\leftarrow(1-t_{b})\,T(x_{\mathrm{data},b})+t_{b}\,x_{\mathrm{data},b}$ $\triangleright$ Interpolation along geodesics
+8:  $\operatorname{\mathcal{L}_{\mathrm{OT}}}(\theta)\leftarrow\sum_{b=1}^{B}\|\nabla_{x}V_{\theta}(x_{t_{b}})+x_{\mathrm{data},b}-T(x_{\mathrm{data},b})\|^{2}$ $\triangleright$ Loss function
+9:  $\theta\leftarrow\theta-\alpha\nabla_{\theta}\operatorname{\mathcal{L}_{\mathrm{OT}}}(\theta)$ $\triangleright$ Gradient update with learning rate $\alpha$
+10:end for
+11:return $\theta\triangleright$ Trained $\theta$
 2.1.2 Contrastive objective $\operatorname{\mathcal{L}_{\mathrm{CD}}}$
@@
 The $\mathrm{sg}(\cdot)$ indicates a stop-gradient operator, which ensures gradients do not back-propagate through the sampling procedure.
-2.1.3 Dual objective
-To balance the flow-like loss $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ with the contrastive divergence loss $\operatorname{\mathcal{L}_{\mathrm{CD}}}$, we adopt the linear temperature schedule (3).
-We define a sampling time $\operatorname{\tau_{s}}$.
-In fact, although convergence to the equilibrium distribution is guaranteed only as $\operatorname{\tau_{s}}\to\infty$, we empirically observe that sample quality (measured with Fréchet inception distance (FID)) plateaus by $\operatorname{\tau_{s}}=3.0$ on CIFAR-10; see Section A.2.
-The time discretization determines the temperature samples $\varepsilon^{(m)}\vcentcolon=\varepsilon(m\Delta t)$.
-We define a dataset-specific hyperparameter $\lambda_{\mathrm{CD}}$ to scale $\operatorname{\mathcal{L}_{\mathrm{CD}}}$.
-The resulting algorithm is described in detail in Algorithm 1 and Algorithm 2.
-Since Algorithm 2 benefits from high-quality negatives, we begin with Algorithm 1 (and, thus, with $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ only) to ensure sufficient mixing of noise-initialized negatives.
-Section A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases.
-Hyper-parameters for each dataset are defined in Appendix D.
+2.1.3 Dual objective and implementation notes
+To balance the deterministic flow-like regime (where $\varepsilon\approx 0$) away from the data manifold and the stochastic Boltzmann regime (where $\varepsilon\approx\varepsilon_{\max}$) near equilibrium, we adopt the linear temperature schedule described in (3).
+We introduce a dataset-specific hyperparameter $\lambda_{\mathrm{CD}}$ to stabilize the contrastive objective by appropriately weighting $\operatorname{\mathcal{L}_{\mathrm{CD}}}$ relative to $\operatorname{\mathcal{L}_{\mathrm{OT}}}$.
+The resulting algorithm is described in detail in Algorithm˜1 and Algorithm˜2.
+Since Algorithm˜2 benefits from high-quality negatives, we begin with Algorithm˜1 (and, thus, with $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ only) to ensure sufficient mixing of noise-initialized negatives.
+Given the trained models, we define a sampling time $\operatorname{\tau_{s}}$.
+Although convergence to the equilibrium distribution is guaranteed only as $\operatorname{\tau_{s}}\to\infty$, we empirically observe that sample quality (measured with Fréchet inception distance (FID)) plateaus by $\operatorname{\tau_{s}}=3.25$ on CIFAR-10; see Section˜A.2.
+The sampling procedure, which optionally includes conditional and interaction terms, is detailed in Algorithm˜3.
+In practice, we implement training using explicit Euler–Maruyama updates and sampling with an Euler–Heun predictor-corrector scheme, while for simplicity the algorithms illustrate only explicit updates.
+Additionally, the constant factor $1/\varepsilon_{\max}$ in $\operatorname{\mathcal{L}_{\mathrm{CD}}}$ is absorbed into $\lambda_{\mathrm{CD}}$.
+Section˜A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases.
+Hyperparameters for each dataset, along with intuitions to guide their selection for new datasets, are provided in Appendix˜D.
 Algorithm 2 Phase 2 (main training).
-1:$\theta\leftarrow\theta_{\text{pretrained}}$triangleright Initialize from the warmed-up model (Algorithm 1)
-2:for iteration $0,1,\ldots$ do
-3:    Sample mini-batch $\{x_{\mathrm{data}}^{(j)}\}_{j=1}^{B}\sim\mathcal{D}$triangleright Mini-batch of data samples
-4:    $\operatorname{\mathcal{L}_{\mathrm{OT}}}\leftarrow$ Use lines 4–7 from Algorithm 1
-5:    Initialize $\{x_{\mathrm{neg}}^{(0)}\}$ from noise and/or data $\triangleright$ Negative samples
-6:    for $m=0,1,\ldots,M_{\text{Langevin}}$ do
-7:         Sample $\eta\sim\mathcal{N}(0,I)$
-8:         $x_{\mathrm{neg}}^{(m+1)}\leftarrow x_{\mathrm{neg}}^{(m)}-\Delta t\nabla_{x}V_{\mathrm{sg}(\theta)}(x_{\mathrm{neg}}^{(m)})+\sqrt{2\Delta t\varepsilon^{(m)}}\eta$triangleright Langevin dynamics given $\varepsilon(t)$
-9:         if $x_{\mathrm{neg}}^{(0)}$ initialized from data then
-10:             Use $\varepsilon_{\max}$ instead of $\varepsilon^{(m)}$triangleright Use maximum temperature if initialized at the data manifold
-11:         end if
-12:    end for
-13:    $\mathcal{L}_{\mathrm{neg}}\leftarrow\mathbb{E}[V_{\theta}(x_{\mathrm{neg}}^{(M)})]$
-14:    $\mathcal{L}_{\mathrm{pos}}\leftarrow\mathbb{E}_{\mathrm{data}}[V_{\theta}(x_{\mathrm{data}})]$$\triangleright$ Compute expected potential energy on the data
-15:    $\operatorname{\mathcal{L}_{\mathrm{CD}}}\leftarrow\mathcal{L}_{\mathrm{pos}}-\mathcal{L}_{\mathrm{neg}}$
-16:    $\mathcal{L}(\theta)\leftarrow\operatorname{\mathcal{L}_{\mathrm{OT}}}+\lambda_{\mathrm{CD}}\operatorname{\mathcal{L}_{\mathrm{CD}}}$
-17:    Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta)$triangleright Gradient descent step
-18:end for
-19:return $\theta$triangleright Trained $\theta$
+[…]
+2:for iteration $n=0,1,\ldots$ do
+3:  $\operatorname{\mathcal{L}_{\mathrm{OT}}}\leftarrow$ Use lines 3–8 from Algorithm 1
+4:  Initialize negative samples $\{x_{\mathrm{neg},b}^{(0)}\}_{b=1}^{B}$ from noise and/or data $\triangleright$ Negative samples
+5:  for $m=0,1,\ldots,M_{\text{Langevin}}-1$ do
+6:    for $b=1$ to $B$ do
+[…]
+8:     Sample $\eta_{b}\sim\mathcal{N}(0,I)$
+9:     $x_{\mathrm{neg},b}^{(m+1)}\leftarrow x_{\mathrm{neg},b}^{(m)}-\Delta t\nabla_{x}V_{\mathrm{sg}(\theta)}(x_{\mathrm{neg},b}^{(m)})+\sqrt{2\Delta t\varepsilon^{(m)}}\,\eta_{b}$ $\triangleright$ Langevin dynamics step
+10:    end for
+11:  end for
+12:  $\operatorname{\mathcal{L}_{\mathrm{CD}}}\leftarrow\frac{1}{B}\sum_{b=1}^{B}\left[V_{\theta}(x_{\mathrm{data},b})-V_{\theta}(x_{\mathrm{neg},b}^{(M_{\text{Langevin}})})\right]$ $\triangleright$ Contrastive divergence loss
+13:  $\mathcal{L}(\theta)\leftarrow\operatorname{\mathcal{L}_{\mathrm{OT}}}+\lambda_{\mathrm{CD}}\operatorname{\mathcal{L}_{\mathrm{CD}}}$
+14:  Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta)$ $\triangleright$ Gradient descent step
+15:end for
+16:return $\theta$ $\triangleright$ Trained $\theta$
 3 Applications
 In this section, we demonstrate the effectiveness and versatility of our proposed Energy Matching approach across three applications: (i) unconditional generation (ii) inverse problems, and (iii) LID estimation.
-The model architecture and all the training details are reported in Appendix D.
+The model architecture and all the training details are reported in Appendix˜D.
 3.1 Unconditional generation
@@
 Our approach, Energy Matching, offers a simple (a single time-independent scalar field) yet powerful EBM-based framework.
-We evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009) and ImageNet32x32 (Deng et al., 2009; Chrabaszcz et al., 2017) datasets, reporting FID scores in Table 1 and Table 2.
-Our method surpasses state-of-the-art EBMs, achieving a $50\%$ reduction in FID.
+We evaluate our approach on CIFAR-10 (Krizhevsky and Hinton, 2009) and ImageNet32x32 (Deng et al., 2009; Chrabaszcz et al., 2017) datasets, reporting FID scores in Table˜1 and Table˜2, respectively.
+Our method outperforms state-of-the-art EBMs, reducing the FID score by more than $50\%$.
 […]
 .
-Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories Ensembles: Diffusion + (one or many) EBMs Diffusion Models Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.54 EGC (Guo et al., 2023) 5.36 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 3.45 Cooperative DRL (40M params) (Zhu et al., 2024) 4.31 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45 Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68 Energy-based Models Flow-based Models ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07 CLEL-large (32M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35 Energy Matching (50M params, Ours) 3.97 OT-CFM (37M params) (Tong et al., 2023) 4.11
+Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories Ensembles: Diffusion + (one or many) EBMs Diffusion Models Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.45 EGC (Guo et al., 2023) 5.36 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 3.45 Cooperative DRL (40M params) (Zhu et al., 2024) 4.31 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45 Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68 Energy-based Models Flow-based Models ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07 CLEL-large (32M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35 Energy Matching (50M params, Ours) 3.34 OT-CFM (37M params) (Tong et al., 2023) 4.04
 […]
@@
 CLEL-large (Lee et al., 2023) (32M params) 15.47
-Energy Matching (50M params, Ours) 7.67
+Energy Matching (50M params, Ours) 6.64
 3.2 Inverse problems
@@
 Samples from this posterior can be drawn by starting from a random sample $x^{(0)}\sim\mathcal{N}(0,I)$ and following a Langevin update.
-We detail the algorithm for generating solutions to inverse problems in Algorithm 3 of Appendix B.
+We detail the algorithm for generating solutions to inverse problems in Algorithm˜3 (which also incorporates additional interaction energy $W(x,x^{\prime})$ between generated samples).
 We demonstrate our model’s capabilities qualitatively through a controlled inpainting task and quantitatively via a protein inverse design benchmark.
+Specific hyperparameters are detailed in Appendix˜D.
+Algorithm 3 Unconditional/conditional sampling with optional interaction energy
+1:for $m=1$ to $M$ do
+2:  Initialize $x_{m}^{(0)}$ from noise and/or data $\triangleright$ Initialize each chain
+3:end for
+[…]
+5:for $n=0,1,\ldots,N-1$ do
+6:  for $m=1,2,\ldots,M$ do $\triangleright$ Prior + data fidelity + interaction
+[…]
+8:    $U_{\theta}\bigl(x_{m}^{(n)}\bigr)\leftarrow V_{\theta}\bigl(x_{m}^{(n)}\bigr)+\varepsilon^{(n)}\bigl\lVert y-A\bigl(x_{m}^{(n)}\bigr)\bigr\rVert^{2}/\zeta^{2}+\varepsilon^{(n)}\sum_{k\neq m}W\bigl(x_{m}^{(n)},x_{k}^{(n)}\bigr)$
+9:    Sample $\eta_{m}^{(n)}\sim\mathcal{N}(0,I)$
+10:    $x_{m}^{(n+1)}\leftarrow x_{m}^{(n)}-\Delta t\,\nabla_{x}U_{\theta}\bigl(x_{m}^{(n)}\bigr)+\sqrt{2\varepsilon^{(n)}\Delta t}\;\eta_{m}^{(n)}$ $\triangleright$ Langevin dynamics step
+11:  end for
+12:end for
+13:return $\{x_{m}^{(N)}\}_{m=1}^{M}$ $\triangleright$ Final samples
 Controlled inpainting.
-Suppose we want to recover two images from a masked image while encouraging diverse reconstructions.
-EBMs allow this by introducing an additional interaction energy, $W(x_{1},x_{2})=-\frac{\|B(x_{1}-x_{2})\|^{2}}{\sigma^{2}}$, where $B$ has ones in the region of interest (focusing diversity there) and zeros elsewhere, and $\sigma$ is a hyper-parameter controlling the interaction’s strength.
-Specifically, we define $p(x_{1},x_{2}\mid y)\propto p(x_{1}\mid y)\,p(x_{2}\mid y)\,\exp\bigl(-W(x_{1},x_{2})\bigr{missing})$, which gives high probability to pairs $(x_{1},x_{2})$ that lie far apart in the specified region $B$.
-This encourages exploring the edges of the posterior rather than just its modes, and with suitable $W$, samples shift toward rare events without needing many draws.
-To illustrate the interaction term’s advantages for diverse reconstruction, we apply our method to a CelebA (Liu et al., 2015) $64\times 64$ inpainting task.
-As shown in Figure 2, we start from a partially observed (masked) face and aim to reconstruct two distinct high-fidelity completions.
+[…]
 […]
@@
 On the right are two reconstructions: the top pair without the interaction term and the bottom pair with it.
-The interaction term applies in the solid red square (where $B$ has ones), and the measurement matrix $A$ is the dotted blue square (zeros inside, ones outside).
+The interaction term applies in the solid red square (where $B$ has ones), and the measurement matrix AA is the dotted blue square (zeros inside, ones outside).
 By encouraging $x_{1}$ and $x_{2}$ to differ in the target region, the interaction yields a wider range of completions while preserving fidelity.
 Protein inverse design.
-In Figure 3, we demonstrate our method’s performance on the inverse design problem of generating Adeno-Associated Virus (AAV) capsid protein segments (Bryant et al., 2021).
+In Figure˜3, we demonstrate our method’s performance on the inverse design problem of generating Adeno-Associated Virus (AAV) capsid protein segments (Bryant et al., 2021).
 Given a desired functional property (fitness)—here defined as the predicted viral packaging efficiency normalized between 0 and 1—the goal is to design novel protein sequences satisfying this target condition.
@@
 This enables a flexible trade-off between fitness and diversity, resulting in high fitness scores alongside substantially improved sequence diversity.
-See Section B.2 for experimental details and dataset descriptions.
+See Appendix˜B for experimental details and dataset descriptions.
 […]
@@
 Estimating the LID at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities.
-We defer the precise definition to Appendix C.
+We defer the precise definition to Appendix˜C.
 Diffusion-based approaches.
-Recent work leverages pre-trained diffusion models to estimate the LID (Kamkari et al., 2024; Stanczuk et al., 2024) by examining the learned score function.
-However, because these models do not learn the score at $t=0$ (the data manifold), their estimates become unreliable there.
-As a result, current methods rely on approximations, for instance, by evaluating the score at small but nonzero times $t=t_{0}$, where the computations remain sufficiently reliable.
+Recent work leverages pretrained diffusion models to estimate the LID (Kamkari et al., 2024; Stanczuk et al., 2024) by examining the learned score function.
+However, since these models do not learn the score at the data manifold ($t=1$), their estimates become unreliable there.
+Consequently, current methods rely on approximations, for instance by evaluating the score in the proximity of the data manifold ($t=1-t_{0}$), where computations remain sufficiently reliable.
 Hessian-based LID Estimation.
@@
 To this end, we compute the Hessian matrix $\nabla_{x}^{2}V(x_{\mathrm{data}})$ at a given data point and perform its spectral decomposition.
-We define near-zero eigenvalues as those whose absolute values lie within a small threshold $\tau$ (in our experiments, we set $\tau=10$ for MNIST (Deng, 2012) and $\tau=2$ for CIFAR-10).
+We define near-zero eigenvalues as those whose absolute values lie within a small threshold $\tau$ (in our experiments, we set $\tau=3$ for MNIST (Deng, 2012) and $\tau=2$ for CIFAR-10).
 The count of near-zero eigenvalues reflects the number of flat directions and thus reveals the local dimension.
-As shown in Table 3, the LID estimates we obtain exhibit stronger correlations with PNG compression size [^2] PNG is a lossless compression scheme specialized for images and can provide useful guidance when no LID ground truth is available Kamkari et al.
+As shown in Table˜3, the LID estimates we obtain exhibit stronger correlations with PNG compression size [^2] PNG is a lossless compression scheme specialized for images and can provide useful guidance when no LID ground truth is available Kamkari et al.
 (2024).
 (evaluated on 4096 images) using Spearman’s correlation.
-Figure 4 offers qualitative illustrations.
+Figure˜4 offers qualitative illustrations.
 Our EBM-based approach compares favorably to diffusion-based methods, as it relies on fewer approximations by performing computations exactly on the data manifold rather than merely in its vicinity.
@@
 NB (Stanczuk et al., 2024) 0.864 0.894
-Energy Matching (Ours) 0.895 0.901
+Energy Matching (Ours) 0.877 0.901
 Table 3: Spearman’s correlation coefficients of LID estimates with PNG compression rate. Benchmarks results reported in (Kamkari et al., 2024).
@@
 Limitations.
-First, our method requires the extra gradient computation with respect to input, which can increase GPU memory usage (e.g.
-20-40%), in particular during training.
-Second, for very high-resolution datasets, computing the full Hessian spectrum ($O(N^{3})$) may be impractical; and partial spectrum methods like random projections or iterative solvers may be used instead.
+First, our method requires an additional gradient computation with respect to the input, which can increase GPU memory usage (e.g., by 20–40%), particularly during training.
+Second, when estimating the LID (Section 3.3) for very high-dimensional datasets, computing the full Hessian spectrum may be impractical due to its computational complexity of $O(d^{3})$; in such cases, partial-spectrum methods such as random projections or iterative solvers can be employed instead.
 Outlook.
-In contrast to the widespread belief, we demonstrated that time-independent curl-free methods for generative flows work and offer an exciting direction for future research.
+Contrary to widespread belief, we demonstrated that time-independent irrotational methods for generative flows are highly effective and offer an exciting direction for future research.
 Our Energy Matching approach has the potential to yield novel insights into controlled generation and inverse problems for cancer research Weidner et al.
@@
 (2024); Molinaro et al.
-(2024) and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.
+(2024), and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.
+Moreover, Energy Matching aligns naturally with recent generative AI trends toward scaling inference for new capabilities (Zhang et al., 2025; Ma et al., 2025), further broadening its potential impact across scientific and engineering domains.
+Acknowledgments and Disclosure of Funding
+This research was supported by the Helmut Horten Foundation and the European Cooperation in Science and Technology (COST).
 Appendix A Additional details on Energy Matching
@@
 A.1 Energy landscape during training
-In Figure 5, we visualize how the potential $V_{\theta}(x)$ transitions from a flow-like regime, where the OT loss enforces nearly zero curvature away from the data manifold (a), to an EBM-like regime, where the curvature around the new data geometry (here, two moons) is adaptively increased to approximate $\log p_{\text{data}}(x)$ (b).
+In Figure˜5, we visualize how the potential $V_{\theta}(x)$ transitions from a flow-like regime, where the OT loss enforces nearly zero curvature away from the data manifold (a), to an EBM-like regime, where the curvature around the new data geometry (here, two moons) is adaptively increased to approximate $\log p_{\text{data}}(x)$ (b).
 This two-stage design yields a well-shaped landscape that is both efficient to sample (thanks to a mostly flat potential between clusters) and accurate for density estimation near the data modes.
@@
 Figure 5: Visualization of the energy $V_{\theta}(x)$ landscapes driving the samples from eight Gaussians to two moons.
-See Figure 1 for the 2D perspective.
+See Figure˜1 for the 2D perspective.
 (a) The OT flow loss enforces zero curvature in $V_{\theta}(x)$ along the trajectories to the target.
@@
 Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, producing a stable Boltzmann distribution from which one can sample.
-Figure 6 illustrates the FID as a function of sampling time $\operatorname{\tau_{s}}$ for models trained under these different regimes.
+Figure˜6 illustrates the FID as a function of sampling time $\operatorname{\tau_{s}}$ for models trained under these different regimes.
 In the case of pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$; this occurs because, once the samples move close to the data manifold, there is no Boltzmann-like potential well to keep them from drifting away.
@@
 [2025].
-In Figure 6 we also report results for different values of the temperature-switching parameter $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation 3).
-11112222333300222244446666888810101010Sampling time $\operatorname{\tau_{s}}$ vs sampling time $\operatorname{\tau_{s}}$$\operatorname{\tau^{\ast}}=0.9,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\tau^{\ast}}=1.0,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
+In Figure˜6 we also report results for different values of the temperature-switching parameter $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation˜3).
+1122330224466881010Sampling time $\operatorname{\tau_{s}}$FIDFID vs sampling time $\operatorname{\tau_{s}}$$\operatorname{\tau^{\ast}}=0.9,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\tau^{\ast}}=1.0,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
 Figure 6: CIFAR-10 unconditional generation FID vs.
-sampling time $\operatorname{\tau_{s}}$ when sampling from models trained under different scenarios: pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ and combined $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, with temperature regime switching parameter $\operatorname{\tau^{\ast}}\in\{0.9,1.0\}$.
+sampling time $\operatorname{\tau_{s}}$ when sampling from models trained under different scenarios: pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ and combined $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, with temperature regime switching parameter $\operatorname{\tau^{\ast}}\in\{0.9,1.0\}$ during sampling.
 Lower FID indicates better generative quality.
 All results for Euler-Heun with $\Delta t=0.01$.
-Appendix B Posterior sampling
-This section provides additional algorithmic details and experimental settings for inverse problem sampling tasks presented in the main text.
-B.1 Algorithm description
-The detailed sampling procedure for inverse problems, optionally incorporating an interaction energy term to explicitly control the diversity among generated samples, is provided in Algorithm 3.
-The Langevin dynamics alternates between evaluating a composite potential $U_{\theta}(x)$, composed of a data fidelity term, a learned energy-based prior $V_{\theta}$, and, optionally, an interaction energy $W(x,x^{\prime})$.
-This interaction term can, for instance, encourage diversity among parallel chains.
-Samples are initialized independently from the Gaussian prior and iteratively refined through gradient-based updates.
-Specific hyperparameters, including the discretization step size $\Delta t$, total sampling time $T$, and the interaction energy strength, are detailed in Appendix D.
-Algorithm 3 Sampling for inverse problems (with optional interaction energy)
-1:for $m=1$ to $M$ do
-2:    Initialize $x_{m}^{(0)}\sim\mathcal{N}(0,I)$triangleright Start each chain from Gaussian prior
-3:end for
-4:$N\leftarrow\lfloor\operatorname{\tau_{s}}/\Delta t\rfloor$triangleright Number of Langevin steps for sampling time $\operatorname{\tau_{s}}$
-5:for $n=0$ to $N-1$ do
-6:    for $m=1$ to $M$ do $\triangleright$ Data fidelity + prior + interaction energy
-7:         $U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}\leftarrow\frac{\varepsilon^{(n)}}{\zeta^{2}}\bigl{\lVert}y-A\bigl{(}x_{m}^{(n)}\bigr{)}\bigr{\rVert}^{2}+V_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}-\frac{\varepsilon^{(n)}}{\sigma^{2}}\sum_{k\neq m}W\bigl{(}x_{m}^{(n)},x_{k}^{(n)}\bigr{)}$
-8:         $\eta\sim\mathcal{N}(0,I)$triangleright Gaussian noise for Langevin step
-9:         $x_{m}^{(n+1)}\leftarrow x_{m}^{(n)}-\Delta t\,\nabla_{x}U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\sqrt{2\varepsilon^{(n)}\Delta t}\;\eta$triangleright Langevin step
-10:    end for
-11:end for
-12:return $\{x_{m}^{(N)}\}_{m=1}^{M}$triangleright Final set of samples
-B.2 AAV inverse design protein generation
-We consider protein fitness optimization for the adeno-associated virus (AAV) in the medium and hard data regime as proposed by [Kirjner et al., 2024].
-The setup for the latent encoding as well as the backbone architecture for the learned prior in the latent space follows [Bogensperger et al., 2025].
-Conditional sampling uses classifier guidance via learned predictor networks $g_{\phi}$ or $\tilde{g}_{\phi}$ to guide samples toward high-fitness regions.
-The CNN-based networks predicting fitness for AAV sequences are directly sourced from [Kirjner et al., 2024] and are trained only on the limited training data for the medium and hard setup, respectively.
-For training we follow Algorithm 1 and Algorithm 2.
-We then sample 128 sequences using Algorithm 3 (same batch size for all baselines), the hyperparameters are detailed in Appendix D.
-Generated sequences are evaluated for fitness using the learned oracle from Kirjner et al.
-[2024].
-Further, they are assessed for diversity within the generated sequences and novelty relative to the training set [Jain et al., 2022].
-As shown in Table 4, our method achieves state-of-the-art fitness while boosting diversity, and incorporating interaction energy in Algorithm 3 further increases diversity with only a minimal impact on fitness.
+A.3 Ablations on the OT Solver
+We evaluate the computational overhead and sensitivity to solver choice of the OT solver employed in our experiments.
+Computational overhead.
+In the CIFAR-10 experiment, the OT solver accounts for roughly 1.5% of the training iteration time during Phase 1 Algorithm˜1.
+This overhead decreases to a negligible level (approximately 0.01%) in Phase 2 Algorithm˜2, where computational costs are predominantly dominated by the generation of negative samples.
+Impact of solver accuracy and complexity.
+Let us (over‑idealise) and model a standardised CIFAR‑10 image as a vector $x\in\mathbb{R}^{d}$, drawn from $\mathcal{N}(0,I_{d})$, with dimensionality $d=32\times 32\times 3=3072$, and paired with an independent Gaussian-noise vector $z\sim\mathcal{N}(0,I_{d})$.
+Each coordinate of the difference (zx)(z-x) thus follows $\mathcal{N}(0,2)$, and the squared Euclidean distance distribution is $|z-x|_{2}^{2}\sim 2\chi_{d}^{2}$, which has mean $2d$, standard deviation $\sqrt{8d}$, and relative spread $\frac{\sqrt{8d}}{2d}\approx 0.025$.
+This demonstrates the "thin-shell" phenomenon, implying that all entries of the cost matrix $C_{ij}=|z_{i}-x_{j}|^{2}_{2}$ concentrate around nearly identical values, consistent with the distance-concentration effect observed by [Aggarwal et al., 2001].
+Consequently, the choice among exact linear programming (LP), entropic regularisation (Sinkhorn), or even random matching should yield nearly identical cumulative optimal-transport costs, despite their different complexities: $O(n^{3}\log n)$ for LP, $O(\frac{n^{2}}{\kappa})$ for Sinkhorn (with regularisation strength $\kappa$), and $O(n)$ for random matching.
+Empirical evidence supporting this is summarised below:
+  • Our CIFAR-10 runs: FID degrades from 3.34 (LP) to 3.37 (random).
+  • [Tong et al., 2023] Table 5: FID scores of 4.44 (LP) vs 4.46 (random) at 100 steps.
+  • [Tong et al., 2023] Fig. D.2: solver accuracy saturates beyond batch size 16 in the 2D task.
+  • [Terpin et al., 2024] App. C.2: LP and Sinkhorn methods produce indistinguishable results unless regularisation is extreme.
+We employ the LP approach for robustness, negligible cost, and no additional hyperparameters.
+A.4 Sampling Time Analysis
+Computing the gradient $\nabla_{x}V_{\theta}(x)$ via automatic differentiation (autograd [Ansel et al., 2024]) introduces additional computational overhead compared to directly evaluating $V_{\theta}(x)$.
+Specifically, on the CIFAR-10 network architecture (see Figure˜7), gradient evaluation is approximately $2.15\times$ slower, as it requires both forward and backward passes.
+In contrast, flow matching and diffusion models directly parameterize the velocity field, thus only requiring forward computations during sampling.
+Nevertheless, despite this per-step computational cost, our method achieves competitive overall sampling efficiency due to a reduced number of integration steps needed for high-quality generation.
+As demonstrated in Table˜4, Energy Matching achieves a lower FID (3.34) in 173 seconds per batch, outperforming OT-FM (FID 3.74 in 136 seconds per batch) and DDPM++ (FID 3.45 in 183 seconds per batch).
+Our results thus indicate a favorable balance between computational overhead per step and total sampling runtime.
+Table 4: Comparison of sampling efficiency and quality on CIFAR-10 (batch size 128, NVIDIA R6000 48GB GPU).
+Despite gradient computation overhead ($\nabla_{x}V_{\theta}(x)$ via backward pass), Energy Matching achieves superior FID scores with competitive wall-clock sampling time.
+Method Params Steps Sampling Time [s]↓ FID↓
+Flow-/Diffusion-based Models
+     OT-FM [Tong et al., 2023] 37M 1000 136 3.74
+     DDPM++ [Kim et al., 2021] 62M 1000 183 3.45
+Energy-based Models
+     Energy Matching (Ours) 50M 325 173 3.34
+Appendix B Details on AAV inverse design protein generation
+We optimize protein fitness for adeno-associated virus (AAV) sequences in the medium and hard data regimes proposed by [Kirjner et al., 2024], using latent encodings and backbone architectures from [Bogensperger et al., 2025].
+Conditional sampling employs classifier guidance via learned predictor networks $g_{\phi}$ or $\tilde{g}_{\phi}$ to steer samples toward high-fitness regions.
+The CNN-based fitness predictors from [Kirjner et al., 2024] are trained only on the limited training data for each regime.
+Training follows Algorithm˜1 and Algorithm˜2.
+We sample 128 sequences using Algorithm˜3, keeping the same batch size across all baselines; detailed hyperparameters are given in Appendix˜D.
+Generated sequences are evaluated for fitness using the learned oracle from [Kirjner et al., 2024], and further assessed for both intra-set diversity and novelty relative to the training sequences [Jain et al., 2022].
+Our approach achieves state-of-the-art fitness while improving diversity (see Table˜5).
+Incorporating interaction energy in Algorithm˜3 further enhances diversity with manageable impact on fitness.
 Table 5: AAV optimization results.
@@
 To start, we need to introduce the concept of local mass, defined as
-[…]
+$M(r)\;=\;\int_{B(x_{\text{data}};r)}p(x)\,dx,$
 where $p(x)$ is the local density and $B(x_{\text{data}},r)$ is a ball of radius $r$ around $x_{\text{data}}$, i.e.
@@
 The LID is then given by:
-$\mathrm{LID}(x_{\mathrm{data}})=d-\lim_{r\to 0}\frac{\log\bigl(M(r)\bigr{missing})}{\log(r)}.$
+$\mathrm{LID}(x_{\mathrm{data}})=d-\lim_{r\to 0}\frac{\log\bigl(M(r)\bigr.)}{\log(r)}.$
 Intuitively, $M(r)$ measures how much probability mass is concentrated in a ball of radius $r$ around $x_{\mathrm{data}}$.
@@
 In the context of contrastive divergence, we assume that data points $x_{\mathrm{data}}$ lie in well-like regions [Hyvärinen, 2006], i.e.:
-[…]
-Conceptually, $V(x)$ can be thought of as an energy function; points where $\nabla V(x_{\mathrm{data}})=0$ are near local minima of this energy, and the Hessian $\nabla^{2}V(x_{\mathrm{data}})$ provides information about local curvature (see Figure 4 for a qualitative illustration).
+$\nabla V(x_{\mathrm{data}})\approx 0\quad\text{and}\quad\nabla^{2}V(x_{\mathrm{data}})\text{ is positive semidefinite (or nearly so).}$
+Conceptually, $V(x)$ can be thought of as an energy function; points where $\nabla V(x_{\mathrm{data}})=0$ are near local minima of this energy, and the Hessian $\nabla^{2}V(x_{\mathrm{data}})$ provides information about local curvature (see Figure˜4 for a qualitative illustration).
 Energy-based density.
@@
 Substituting the local quadratic form of $p(x)$ near $x_{\mathrm{data}}$ into the definition of the local mass $M(r)$, we obtain:
-[…]
+$M(r)=\int_{B(x_{\text{data}},r)}p(x)dx\propto\int_{B(x_{\text{data}},r)}\exp\left(-\tfrac{1}{2\varepsilon}(x-x_{\mathrm{data}})^{\top}\nabla_{x}^{2}V(x_{\mathrm{data}})(x-x_{\mathrm{data}})\right)dx.$
 For small $r$, the dominant contribution depends on the rank of the Hessian $\nabla_{x}^{2}V(x_{\mathrm{data}})$.
@@
 Below, we detail the training configurations for CIFAR-10, ImageNet 32x32, CelebA, MNIST, and AAV.
+Additionally, we provide intuitions for practical hyperparameter choices to facilitate effective training across additional datasets.
+We recommend using SiLU activation functions wherever possible, as they smooth out the energy landscape and improve the numerical stability of the $\nabla_{x}V(x)$ computation.
 The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd [Ansel et al., 2024].
-We optimize all models with the Adam optimizer Kingma and Ba [2014], and for CIFAR-10, ImageNet 32x32, and CelebA we maintain an exponential moving average (EMA) of the weights.
+We optimize all models using the Adam optimizer [Kingma and Ba, 2014] and maintain an exponential moving average (EMA) of the model weights.
 While we specifically adopt (i) a one-sided trimmed mean of negative sample energies and (ii) clamping of the contrastive loss for stability, any commonly used EBM technique (e.g., persistent contrastive divergence [Tieleman, 2008], replay buffers, multi-scale negative sampling) could be readily employed.
@@
 […]
-$\beta$ $=\text{clamp threshold for $\mathcal{L}_{\mathrm{CD}}$ (i.e., we clamp $\mathcal{L}_{\mathrm{CD}}$ to be $\geq-\beta$).}$.
-CIFAR-10.
-The architecture is shown in Figure 7.
+[…]
+CIFAR-10:
+The architecture is shown in Figure˜7.
 We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively time-independent) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output.
-Hyperparameters are: $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$.
-We train for 200k iterations using Algorithm 1 with EMA $0.9999$ and then 10k more with Algorithm 2 and EMA $0.999$ on 4xA100.
-The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=8\times 10^{-3},\alpha=0.1,\beta=-0.05$.
-ImageNet 32x32.
-The architecture is shown in Figure 7 (same as for CIFAR-10).
-Hyperparameters are: $\tau_{s}=1.5$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$.
-We train for 600k iterations using Algorithm 1 with EMA $0.9999$ and then 7k more with Algorithm 2 and EMA $0.999$ on 7xA100.
-The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=2\times 10^{-4}.$ $\alpha=0.2,\beta=-0.01$.
-CelebA.
-We scale the CIFAR-10 model by 2$\sim$ 2$\times$ 2 ×; see Figure 8.
-We keep $\tau_{s}=2.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm 1 with EMA $0.9999$ then 7k with Algorithm 2 and EMA $0.999$ on 4xA100.
-The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
-MNIST.
-We train on a single RTX6000 directly with Algorithm 2 for 300k iterations using $\Delta t=0.05$, $M_{\text{Langevin}}=50$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=0.25$.
-The architecture consists of a two-stage CNN followed by a two-layer MLP, comprising a total of 2M parameters.
-AAV.
+Hyperparameters are: $\tau_{s}=3.25$, $\tau^{*}=1.0$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$.
+We train for 145k iterations using Algorithm˜1 with EMA $0.9999$ and then 2k more with Algorithm˜2 and EMA $0.99$ on 4xA100.
+The batch size is 128, learning rate is $1.2\times 10^{-3}$, $\varepsilon_{\max}=0.01$, $\lambda_{\mathrm{CD}}=1\times 10^{-3}$, $\alpha=0.1$, and $\beta=0.02$.
+Negatives initialized on the data manifold follow the same temperature schedule as those initialized from the noise.
+ImageNet 32x32:
+The architecture is shown in Figure˜7 (same as for CIFAR-10).
+Hyperparameters are: $\tau_{s}=2.5$, $\tau^{*}=1.0$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$.
+We train for 640k iterations using Algorithm˜1 with EMA $0.9999$ and then 1k more with Algorithm˜2 and EMA $0.99$ on 7xA100.
+The batch size is 128, learning rate is $6\times 10^{-4}$, $\varepsilon_{\max}=0.01$, $\lambda_{\mathrm{CD}}=1\times 10^{-3}$, $\alpha=0.1$, and $\beta=0.02$.
+Negatives initialized on the data manifold follow the same temperature schedule as those initialized from the noise.
+CelebA:
+We scale the CIFAR-10 model by $\sim 2\times$; see Figure˜8.
+We set $\tau_{s}=2.0$, $\tau^{*}=1.0$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm˜1 with EMA $0.9999$ then 4k with Algorithm˜2 and EMA $0.99$ on 4xA100.
+The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
+MNIST:
+We downscale the CIFAR-10 model (Figure˜7) to 2M parameters by reducing the UNet base width to 32 channels, using channel multipliers [1, 2, 2], setting the number of attention heads in the UNet to 2, simplifying the Transformer head to an embedding dimension of 128, 2 Transformer layers, 2 attention heads, and adjusting the output scale to 100.0.
+We set $\tau_{s}=2.0$, $\tau^{*}=1.0$, $\Delta t=0.025$, $M_{\text{Langevin}}=75$, and train for 50k iterations using Algorithm˜1 with EMA $0.999$ then 3.3k with Algorithm˜2 and EMA $0.99$ on a single A100.
+The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, $\lambda_{\mathrm{CD}}=1\times 10^{-3}$, $\alpha=0.0$, and $\beta=0.05$.
+Negatives initialized on the data manifold follow the same temperature schedule as those initialized from the noise.
+AAV:
 We adopt the one-dimensional CNN architecture as used in [Bogensperger et al., 2025], summing the final-layer activations to obtain the potential.
-We train for 10k iterations using Algorithm 1 and for 1k iterations using Algorithm 2 on a single GeForce RTX 4090.
-The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
-For Algorithm 3 we use $\tau_{s}=1.7$ for AAV medium and $\tau_{s}=1.3$ for AAV hard, $\tau^{*}=0.9$, $\zeta=0.01$ for AAV medium and $\zeta=0.009$ for AAV hard and $\Delta t=0.01$.
+We train for 10k iterations using Algorithm˜1 and for 1k iterations using Algorithm˜2 on a single A100.
+The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, $M_{\text{Langevin}}=200$, $\Delta t=0.01$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
+For Algorithm˜3 we use $\tau_{s}=1.7$ for AAV medium and $\tau_{s}=1.3$ for AAV hard, $\tau^{*}=0.9$, $\zeta=0.01$ for AAV medium and $\zeta=0.009$ for AAV hard.
 We set the target fitness to $y=1$ to aim for the maximum fitness in the generated sequences.
-Input(3$\times$32$\times$32)UNet(37M params)Transformer(PatchEmbed+ViT, 12M params)ScalarPotential
+Intuition for other datasets:
+It is essential for negative samples to reach the equilibrium distribution induced by the model or at least the proximity of the data manifold.
+The condition $M_{\text{Langevin}}\times\Delta t\gg 1$ is critical to achieving this, with $\Delta t$ small enough to ensure that negative samples remain of sufficient quality—typically the same $\Delta t$ as used in flow matching for the given generation task.
+In practice, we set $M_{\text{Langevin}}\times\Delta t=2$ across most experiments.
+We recommend starting with $\tau^{*}=1.0$ to ensure optimal transport regularization near the data manifold, thereby enhancing training stability.
+If additional conditions are required during sampling, exploring lower values ($\tau^{*}<1.0$) may be beneficial, as this parameter does not need to remain consistent between training and sampling (as shown in Figure˜6).
+Training with $\tau^{*}<1.0$ is advised only in special cases, such as extremely low-dimensional problems like that shown in Figure˜1, where it is possible to simultaneously be far from the data manifold (from the perspective of the target mode) and close to it (from the perspective of another mode).
+The parameter $\varepsilon_{\max}$ controls how extensively negative samples explore the space.
+For unconditional generation, we use $\varepsilon_{\max}=0.01$, but for inverse problems or design tasks, higher values (e.g., $\varepsilon_{\max}=0.05$) can improve robustness.
+The parameter $\tau_{s}$ significantly depends on the task (unconditional or conditional) and thus must be tuned accordingly post-training.
+Without explicit tuning, selecting $\tau_{s}=M_{\text{Langevin}}\times\Delta t$ is a reasonable default.
+Finally, parameters $\lambda_{\mathrm{CD}}$, $\alpha$, and $\beta$ influence the stability of contrastive training and must be empirically determined alongside appropriate early stopping.
+[…]
 Hyperparameter Value
-Image size 3$\times$32$\times$32
+[…]
 Base channels (UNet) 128
@@
 Output scale 1000.0
-Figure 7: Diagram of our UNet+Transformer EBM for CIFAR-10 and ImageNet 32x32.
-A UNet (37M params) processes a 3$\times$32$\times$32 image; its output is fed into a Transformer head (PatchEmbed + 8-layer ViT, 12M params) that produces a scalar potential.
-Here we employ the identical UNet architecture as in [Tong et al., 2023], but with the time parameter fixed at $t=0$ to render the model time-independent.
-Input(3$\times$64$\times$64)UNet(83M params)Transformer(PatchEmbed+ViT, 25M params)ScalarPotential
+[…]
+[…]
 Hyperparameter Value
-Image size 3$\times$64$\times$64
+[…]
 Base channels (UNet) 128
@@
 Output scale 1000.0
-Figure 8: Diagram of our UNet+Transformer EBM for CelebA.
-A UNet (83M params) processes a 3$\times$64$\times$64 image; its output is fed into a Transformer head (PatchEmbed+8-layer ViT, 25M params) that produces a scalar potential.
+[…]